### v1

# On braided simple extensions and braided non-semisimple near-group categories

Abstract.
Let $\mathcal{C}$ be a finite tensor category with a decomposition $\mathcal{C}\cong\mathcal{D}\oplus{\mathcal{M}}$ into full abelian categories such that $\mathcal{D}$ is a tensor subcategory and $\mathcal{{M}}$ has a single simple projective object. We call $\mathcal{C}$ a simple extension of $\mathcal{D}$. When $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided subcategory, we will call $\mathcal{C}$ a braided simple extension.
Applying the notion of a simple extension to a pointed non-semisimple category $\mathcal{D}$ gives a natural generalization to the family of near-group categories, which was first introduced by [S1]. Braided near-group categories were fully classified in [S2] and [T].
In this paper, we study the structure of braided simple extensions of a braided finite tensor category $\mathcal{D}$ in general, and specifically when $\mathcal{D}$ is pointed. In particular, we classify non-degenerate braided non-semisimple near-group categories, and prove that any braided non-semisimple near-group category $\mathcal{C}$ is ”an extension” of a non-degenerate braided non-semisimple near-group category by $\text{Rep}(G)$, where $G$ is the Picard group of a ”canonical” symmetric subcategory of $\mathcal{C}$.
1. Introduction
Let $\mathcal{D}$ be a finite tensor category; that is, an abelian, $\mathbbm{k}$-linear, rigid monoidal category with a finite number of (isomorphism classes of) simple objects, where $\mathbbm{1}$ is a simple object and each object has a projective cover. For a monoidal category $\mathcal{C}$, a category $\mathcal{M}$ is called a left $\mathcal{C}$-module category if there exists a bifunctor $\tilde{\otimes}:\mathcal{C}\times\mathcal{M}\to\mathcal{M}$ and a natural isomorphism
$m_{X,Y,M}:(X\otimes Y)\tilde{\otimes}M\to X\tilde{\otimes}(Y\tilde{\otimes}M)$
for $X,Y\in\mathcal{C}$ and $M\in\mathcal{M}$ which is compatible with the associativity constraint of $\mathcal{C}$ and $M\mapsto\mathbbm{1}\tilde{\otimes}M$ is an autoequivalence of $\mathcal{M}$.
A classical result in the theory of finite tensor categories is that if $\mathcal{D}$ has a fiber functor, then Vec is a semisimple $\mathcal{D}$-bimodule category, and $\mathcal{C}:=\mathcal{D}\oplus\text{Vec}$ is also a $\mathcal{D}$-bimodule category (see [EGNO]). A natural question that arises is: when can we define a monoidal structure on $\mathcal{C}$. To be more precise, we ask:
Question 1.0.1.
For a finite tensor category $\mathcal{D}$ with a fiber functor $F:\mathcal{D}\to\text{Vec}$, is there a category $\mathcal{M}$ which is equivalent to Vec as a $\mathcal{D}$-bimodule category, such that $\mathcal{C}:=\mathcal{D}\oplus\mathcal{M}$ has a structure of a finite tensor category (which is compatible with the $\mathcal{D}$-action on $\mathcal{M}$)?
When the answer to this question is yes, we say that $(\mathcal{C},\mathcal{D},\mathcal{M})$ is a simple extension of $\mathcal{D}$. Usually, we will omit $\mathcal{M}$ and simply say that $\mathcal{C}$ is a simple extension of $\mathcal{D}$. Notice that, by definition, if $\mathcal{C}$ is a simple extension of $\mathcal{D}$, then $\mathcal{C}$ has a simple projective object $Q$, the unique simple object in $\mathcal{M}$.
In fusion categories, there is a family of categories called near-group categories, which were introduced in [S1] (though classification of such categories appeared even earlier in [TY]). A near-group category is a fusion category $\mathcal{C}$ with a unique non-invertible object $Q$. Denoting the group of simple invertible objects of $\mathcal{C}$ by $G=\{L_{i}\}$, then a key parameter in the structure of $\mathcal{C}$ is the parameter $r\in\mathbb{N}$ such that $Q\otimes Q\cong\oplus_{i}L_{i}\oplus rQ$. This near-group category is usually denoted by $\mathcal{C}(G,r)$, and we say that $\mathcal{C}$ has a fusion rule $(G,r)$.
There are many known results regarding near-group categories. For instance, in their work, [TY] proved that every near-group category with fusion rule $(G,0)$ must be non-strict, that is, must have non-trivial associativity constraint (they also determine the structure constants of the possible associativity constraints in such categories). In addition, [S2] classified braided near-group categories with fusion rule $(G,0)$, and their braid structures (including the structure of the underlying group $G$), and [T] classified braided near-group categories with fusion rule $(G,r)$ for $r>0$. These results strongly rely on two factors:
  • The semisimplicity of near-group categories, which enables translating the diagrams defining (braided) fusion categories to equations describing the fusion rules $(G,r)$, due to their unique structure (as a fusion category with a unique non-invertible object).
  • The classification of pointed fusion categories, as categories of the form $\text{Vec}^{\omega}_{G}$ of $G$-graded vector spaces, with associativity constraint corresponding to a $3$-cocycle $\omega\in H^{3}(G,\mathbbm{k}^{*})$.
A natural way to generalize near-group categories to the class of finite tensor categories is:
Definition 1.0.2.
A non-semisimple finite tensor category $\mathcal{C}$ is called a non-semisimple near-group category if $\mathcal{C}$ is a simple extension of a non-semisimple pointed finite tensor category $\mathcal{D}$.
Extending the notion from fusion categories, we say that a non-semisimple near-group category $\mathcal{C}$ has generalized fusion rule $(G,r)$ if $G$ is the group of invertible objects of $\mathcal{C}$ (that is, the Picard group of $\mathcal{D}$, of which $\mathcal{C}$ is its simple extension) and $r$ is the number of copies of the unique simple projective object $Q$ in the tensor product $Q\otimes Q$.
[…]
In this paper, we wish to take a first step toward classifing non-semi-simple near-group categories, by focusing on braided non-semisimple near-group categories, i.e. braided simple extensions of braided pointed finite tensor categories, admitting a fiber functor. Such a classification should be easier, due to the classification of braided pointed finite tensor categories with a fiber functor, which is given in [BN1].
Formally, in this work we consider triplets of simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$ for which $\mathcal{D}$ is braided, $\mathcal{M}$ is equivalent as a $\mathcal{D}$-bimodule category to Vec and $\mathcal{C}$ is the simple extension of $\mathcal{D}$, such that $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided tensor subcategory of $\mathcal{C}$. An example for such a category is given in [GS]. On the other hand, there are braided finite tensor categories with simple extensions that are not braided - for instance $\text{Rep}(H_{4})$ is braided but $\text{Rep}(A^{\prime\prime}_{C_{4}})$ is not, because $A^{\prime\prime}_{C_{4}}$ has no quasi-triangular structure (see [M]).
In the fusion case, [T] proves that there are braided near-group categories with fusion rule $(G,r)$ for $r>0$. In this paper, we prove that this fails in non-semisimple near-group categories. That is, we prove the following theorem:
Theorem 1.
Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$. In particular, a non-semisimple braided near-group category is weakly integral.
In their work ([BN2], [BN3]), the authors defined the notion of a central exact sequence of finite tensor categories as an exact sequence $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ of finite tensor categories with exact tensor functors $i:\mathcal{E}\to\mathcal{C}$ and $F:\mathcal{C}\to\mathcal{D}$ such that $i$ is a full embedding, $F$ is surjective and normal, meaning that for every $Y\in\mathcal{C}$ there exists a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the maximal subobject of $F(X)$ that is a direct sum of copies of $\mathbbm{1}_{\mathcal{D}}$, and $F\circ i$ is a fiber functor of $\mathcal{E}$. Such an exact sequence is called braided if the categories $\mathcal{E},\mathcal{C},\mathcal{D}$ and the functors $i,F$ are braided, and it is called a modularization exact sequence if $\mathcal{D}$ is modular (i.e, non-degenerate). In this paper, we prove that braided non-semisimple near-group categories are ”extensions” of non-degenerate braided non-semisimple near-group categories by symmetric fusion categories. Formally, we prove:
Theorem 2.
Let $\mathcal{C}$ be a braided non-semisimple near-group category, then there exist a non-degenerate braided non-semisimple near-group category $\mathcal{D}$ and a finite group $G$ such that $\mathcal{C}$ fits in a modularization exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}.$
The proof of this theorem is constructive, and $(\mathcal{D},G)$ are determined by $\mathcal{C}$. As a result of this theorem, classifying braided non-semisimple near-group categories restricts to classification of non-degenerate braided non-semisimple near-group categories.
In their work, [GS] classified non-degenerate braided finite tensor categories with a Lagrangian subcategory braid equivalent to $\textrm{sRep}(W)$ (the category of supermodules of the exterior algebra over a purely odd supervector space $W$), and they proved that some of these categories are simple braided extensions of $\textrm{sRep}(W\oplus W^{*})$ (though they do not use the term ”braided simple extension”). In this paper, we prove that these examples are, in a way, unique. To be exact, we prove:
Theorem 3.
Every non-degenerate non-semisimple braided near-group category $C$ is a braided simple extension of a category $\mathcal{D}$ braid equivalent to $\textrm{sRep}(W\oplus W^{*})$, with $\textrm{sRep}(W)$ a Lagrangian subcategory, for some purely odd supervector space $W$.
Notice that $\mathcal{D}$ is equivalent to $\textrm{sRep}(W\oplus W^{*})$, with a non-symmetric braid structure (as the theorem states, $\textrm{sRep}(W)$ is the maximal symmetric subcategory of $\mathcal{D}$, i.e. the maximal subcategory in which the restriction of the braiding is symmetric).
The combination of 2 and 3 gives a complete description of the class of braided non-semisimple near-group categories. This classification give two results:
Corollary 4.
Every braided non-semisimple near-group category is not integral.
Theorem 5.
Every braided non-semisimple near-group category $\mathcal{C}$ corresponds to a pair $(G,n_{p})$ such that:
  1. (1)
    $G$ is a group of order $2^{n_{g}}$ with a central element of order $2$.
  2. (2)
    $n_{p}\in\mathbb{N}$ and for $P_{\mathbbm{1}}$ the projective cover of $\mathbbm{1}_{\mathcal{C}}$, $\dim(P_{e})=2^{n_{p}}$ as a vector space.
  3. (3)
    $n_{p}+n_{g}$ is an odd number.
This result is a slightly weaker version of the result given in [S2] concerning braided (fusion) near-group categories. In particular, this Theorem gives a categorical proof to the result of [M] stated above that there is no braided simple extension to $\text{Rep}(H_{4})$, in $\text{Rep}(H_{4})$ we have $n_{g}+n_{p}=2$.
1.1. Structure of the paper
In Section 2, we give the necessary preliminaries, including results concerning the Müger centralizer in braided finite tensor categories, which will be used extensively.
In Section 3 we consider general braided simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$, and develop some basic tools that will be required to the study of braided non-semisimple near-group categories. In Section 4 we focus on braided non-semisimple near-group categories.
Acknowledgments
I’m grateful to Inna Entova-Aizenbud for helpful discussions. This research was supported by the Israel Science Foundation grant ISF No. 1362/23 (PI: Inna Entova-Aizenbud).
2. Preliminaries
In this section, we recall some basic background on finite tensor categories and braiding of tensor categories. Throughout this paper, we work over $\mathbbm{k}$, an algebraically closed field of characteristic zero, and all categories are assumed to be abelian and $\mathbbm{k}$-linear.
2.1. Notation and terminology
By a finite tensor category $\mathcal{C}$, we mean a $\mathbbm{k}$-linear, abelian, monoidal category with finitely many simple objects (up to isomorphism), such that each simple object has a projective cover, and $\mathbbm{1}$ is a simple object. Throughout this paper, we assume all the tensor categories have trivial unit constraints, but not necessarily trivial associativity constraint. $\mathcal{C}$ is called semisimple if every object is a direct sum of simple objects.
For an abelian category $\mathcal{C}$, a full subcategory $\mathcal{D}\subset\mathcal{C}$ is called a Serre subcategory if for every short exact sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{C}$, $Y\in\mathcal{D}$ if and only if $X,Z\in\mathcal{D}$. In particular, a Serre subcategory is closed under taking subobjects, quotients and extensions.
In a finite tensor category, every object $X$ has a socle filtration $0=X_{0}\subset X_{1}\subset...\subset X_{n}=X$ in which $X_{i+1}/X_{i}$ is the sum of simple subobjects of $X/X_{i}$. For $X\in\mathcal{C}$, we define the Loewy length to be the length of the socle filtration of $X$.
A tensor functor $F:\mathcal{C}\to\mathcal{D}$ is called surjective if every $Y\in\mathcal{D}$ is a subquotient of $F(X)$ for some $X\in\mathcal{C}$. Given a finite tensor category $\mathcal{C}$, a fiber functor is an exact faithful functor $F:\mathcal{C}\to\text{Vec}$ such that $F(\mathbbm{1})=\mathbbm{k}$ with a family of isomorphisms $\{J_{X,Y}:F(X)\otimes F(Y)\to F(X\otimes Y)|X,Y\in\mathcal{C}\}$ which satisfy the axioms of a monoidal functor.
By definition, a finite tensor category $\mathcal{D}$ has a simple extension $\mathcal{C}$ if $\mathcal{D}$ has a bimodule category which is semisimple with a unique simple object, in other words, if $\mathcal{D}$ has a fiber functor. This implies that $\mathcal{D}$ is equivalent (as a monoidal category) to a category of representations of some Hopf algebra, thus has trivial associativity constraint.
2.2. Braiding, symmetry, twists and Müger centralizer
A tensor category $\mathcal{C}$ is called braided if there exist isomorphisms $\{c_{X,Y}:X\otimes Y\to Y\otimes X\}_{X,Y\in\mathcal{C}}$ such that the following holds for objects $X,Y,Z\in\mathcal{C}$ and morphisms $\psi:X\to Z$:
(1) $c_{X\otimes Y,Z}=a_{Z,X,Y}\circ(c_{X,Z}\otimes\text{id}_{Y})\circ a^{-1}_{X,Z,Y}\circ(\text{id}_{X}\otimes c_{Y,Z})\circ a_{X,Y,Z},$
(2) $(\text{id}_{Y}\otimes\psi)\circ c_{X,Y}=c_{Z,Y}\circ(\psi\otimes\text{id}_{Y})$
for $a$ the associativity constraint of $\mathcal{C}$. For more details, see for instance [EGNO].
For a braided finite tensor category $\mathcal{C}$, let $s$ be the squared braiding, i.e., $s_{X,Y}:=c_{Y,X}\circ c_{X,Y}$ for every $X,Y\in\mathcal{C}$. Recall that two objects $X,Y\in{\mathcal{C}}$ centralize each other if $s_{X,Y}=\text{id}_{X\otimes{Y}}$, and that the centralizer $\mathcal{D}^{\prime}$ of a full tensor subcategory $\mathcal{D}\subseteq\mathcal{C}$ is the full subcategory of $\mathcal{C}$ consisting of all objects which centralize every object of $\mathcal{D}$ (see, e.g., [DGNO]). The notion of Müger’s centralizer enables us to distinguish between different types of braided finite tensor categories. For finite braided tensor categories $\mathcal{D}\subset\mathcal{C}$ we say that:
  • $\mathcal{D}$ is symmetric if $D\subseteq\mathcal{D}^{\prime}$.
  • $\mathcal{D}$ is Lagrangian if $\mathcal{D}=\mathcal{D}^{\prime}$.
  • $\mathcal{D}$ is central in $\mathcal{C}$ if $\mathcal{D}^{\prime}=\mathcal{C}$ (thus a central category is always symmetric).
  • $\mathcal{C}$ is non-degenerate if $\mathcal{C}^{\prime}=\text{Vec}$ (equivalently, Vec is a Lagrangian subcategory of $\mathcal{C}$).
In general, for $X,Y\in\mathcal{C}$ centralizing each other we have $s_{X,Y}=\text{id}_{X\otimes Y}\Leftrightarrow s_{Y,X}=\text{id}_{Y\otimes X}$.
By [Sh], we have in general for $\mathcal{D}\subset\mathcal{C}$
(3) $\text{\rm FPdim}(\mathcal{D})\text{\rm FPdim}(\mathcal{D}^{\prime})=\text{\rm FPdim}(\mathcal{C})\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$
When considering $\mathcal{D}\subset\mathcal{C}$, the centralizer of $\mathcal{D}$ in $\mathcal{D}$ is not necessarily its centralizer in $\mathcal{C}$. When needed, to avoid confusion, we will denote the centralizer of $\mathcal{D}$ in $\mathcal{D}$ by $(\mathcal{D})^{\prime}_{\mathcal{D}}$.
This also applies for three categories $\mathcal{E}\subset\mathcal{D}\subset\mathcal{C}$. Hence $(\mathcal{E})^{\prime}_{\mathcal{D}}$ might differ from $(\mathcal{E})^{\prime}_{\mathcal{C}}$ or $(\mathcal{E})^{\prime}_{\mathcal{E}}$. In all cases, the centralizer in the largest category will always be denoted without the subscript, i.e., in this case, $(\mathcal{E})^{\prime}_{\mathcal{C}}$ will simply be denoted as $\mathcal{E}^{\prime}$. Clearly, if $\mathcal{E}\subset\mathcal{D}\subset\mathcal{C}$ then $\mathcal{E}^{\prime}\subset(\mathcal{E})^{\prime}_{\mathcal{D}}\subset(\mathcal{E})^{\prime}$$_{\mathcal{E}}$.
For any $X,Y,Z\in\mathcal{C}$ we have the following equality, which is a direct calculation using the definition of braiding:
(4) $s_{X\otimes Y,Z}=a^{-1}_{X,Y,Z}\circ(\text{id}_{X}\otimes c_{Z,Y})\circ a_{X,Z,Y}\circ(s_{X,Z}\otimes\text{id}_{Y})\circ a^{-1}_{X,Z,Y}\circ(\text{id}_{X}\otimes c_{Y,Z})\circ a_{X,Y,Z}$
(5) $s_{X,Y\otimes Z}=a_{X,Y,Z}\circ(c_{Y,X}\otimes\text{id}_{Z})\circ a^{-1}_{Y,X,Z}\circ(\text{id}_{Y}\otimes s_{X,Z})\circ a_{Y,X,Z}\circ(c_{X,Y}\otimes\text{id}_{Z})\circ a^{-1}_{X,Y,Z}$
2.3. De-equivariantization
In this subsection, we briefly recall the notion of de-equivariantization and some results which we will use in the classification of braided non-semisimple near-group categories. For more details, see [EGNO].
For a finite tensor category $\mathcal{C}$, and an algebra object $A\in\mathcal{C}$, $\textrm{Mod}_{\mathcal{C}}(A)$ is the category of right $A$-modules in $\mathcal{C}$. When $\mathcal{C}$ is braided and $A\in\mathcal{C}^{\prime}$, the category $\textrm{Mod}_{\mathcal{C}}(A)$ becomes a finite tensor category, with tensor product $\otimes_{A}$, and the free module functor $F:\mathcal{C}\to\textrm{Mod}_{\mathcal{C}}(A)$ given by $F(X)=X\otimes A$ is a surjective tensor functor.
For $X,Y\in\textrm{Mod}_{\mathcal{C}}(A)$, we denote by $\text{Hom}_{A}(X,Y)$ the vector space of $A$-module morphisms $f:X\to Y$ which are morphisms in the category $\mathcal{C}$. Notice that in general, we don’t have $\text{Hom}_{A}(X,Y)=\text{Hom}_{\mathcal{C}}(X,Y)$. For $X\in\mathcal{C}$ and $Y\in\textrm{Mod}_{\mathcal{C}}(A)$ we do have:
$\text{Hom}_{A}(X\otimes A,Y)=\text{Hom}_{\mathcal{C}}(X,\bar{Y})$
for $\bar{Y}$ the image of $Y$ under the forgetful functor $\textrm{Mod}_{\mathcal{C}}(A)\to\mathcal{C}$ (see Lemma 7.8.12 in [EGNO]).
It is well known that for a finite group $G$, $A_{G}:=\text{Fun}(G,\mathbbm{k})$ is an algebra object in the category $\text{Rep}(G)$, see for instance [EGNO].
Proposition 2.3.1.
Let $\mathcal{C}$ be a braided finite tensor category with braiding $c$, $G$ a finite group, and assume that $\text{Rep}(G)\subset\mathcal{C}^{\prime}$ as a symmetric subcategory, then $\textrm{Mod}_{\mathcal{C}}(A_{G})$ is a braided finite tensor category with braiding $\tilde{c}:M\otimes_{A_{G}}N\to N\otimes_{A_{G}}M$ such that the diagram
$M\otimes N$$N\otimes M$$M\otimes_{A_{G}}N$$N\otimes_{A_{G}}M$$\scriptstyle{c_{M,N}}$$\scriptstyle{\tilde{c}_{M,N}}$
commutes, and the free module functor functor $F:\mathcal{C}\to\textrm{Mod}_{\mathcal{C}}(A_{G})$ is braided.
Proof.
The proof is parallel to Proposition 8.23.8 in [EGNO], using the inclusion as a fully faithful braided tensor functor $\text{Rep}(G)\to\mathcal{C}^{\prime}$. ∎
For a braided finite tensor category $\mathcal{C}$ for which 2.3.1 holds, we denote $\textrm{Mod}_{\mathcal{C}}(\text{Fun}(G,\mathbbm{k}))$ as $\mathcal{C}_{G}$, and call this category the $G$-de-equivariantization of $\mathcal{C}$.
Remark 2.3.2.
In the case of fusion categories, if $\mathcal{C}$ has a $G$-de-equivariantization, then we have $(\mathcal{C}_{G})^{G}\simeq\mathcal{C}$, for $\mathcal{D}^{G}$ the $G$-equivariantization of $\mathcal{D}$ by $G$ (see [EGNO] for details on $G$-equivariantization). In finite tensor categories, this is not always true. For example, for $H$ the Sweedler’s Hopf algebra, $\text{Rep}(\mathbb{Z}_{2})$ is central in $\text{Rep}(H)$, and $\text{Rep}(H)_{\mathbb{Z}_{2}}=\text{Vec}$, so $(\text{Rep}(H)_{\mathbb{Z}_{2}})^{\mathbb{Z}_{2}}\not\simeq\text{Rep}(H)$. This happens because the fiber functor $F:\text{Rep}(H)\to\text{Vec}$ is not a braided fiber functor.
Proposition 2.3.3.
Let $\mathcal{C}$ be a braided finite tensor category with de-equivariantization $\mathcal{C}_{G}$, and $X\in\mathcal{C}$ a simple object. If $X\in\mathcal{C}$ is invertible, then $X\otimes A$ is invertible in $\mathcal{C}_{G}$.
Proof.
The free module functor is a surjective tensor functor, so by [EO] it preserves Frobenious-Perron dimensions, thus $\text{\rm FPdim}_{\mathcal{C}_{G}}(X\otimes A)=1$, hence by [EGNO], $X\otimes A$ is invertible in $\mathcal{C}_{G}$. ∎
2.4. Braided exact sequences of tensor categories
In this subsection, we recall the notion of braided exact sequences of finite tensor categories, which was introduced in [BN3].
A tensor functor $F:\mathcal{C}\to\mathcal{D}$ is called normal if any object $X\in\mathcal{C}$ admits a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the largest subobject of $F(X)$ isomorphic to $(\mathbbm{1}_{\mathcal{D}})^{\oplus n}$ for $n\in\mathbb{N}$.
For $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ finite tensor categories over $\mathbbm{k}$, and tensor functors $i:\mathcal{C}_{1}\to\mathcal{C}_{2}$, $F:\mathcal{C}_{2}\to\mathcal{C}_{3}$, the sequence of tensor categories:
$\mathcal{C}_{1}\xrightarrow{i}\mathcal{C}_{2}\xrightarrow{F}\mathcal{C}_{3}$
is called an exact sequence of tensor categories if $F$ is surjective and normal, and $i$ is a full embedding such that $F\circ i(X)\in\text{Vec}$ for all $X\in\mathcal{C}_{1}$.
Remark 2.4.1.
Notice that this definition requires in particular that $\mathcal{C}_{1}$ admits a fiber functor. The notion of exact sequences was generalized to avoid this restriction by [EG], but in this work, we will consider exact sequences as defined in [BN3].
An exact sequence of tensor categories is called braided if the categories $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ are braided and the functors $i,F$ are braided. A braided exact sequence is called modularization exact sequence if $\mathcal{C}_{3}$ is non-degenerate.
Remark 2.4.2.
If $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ is a braided exact sequence, then $\mathcal{E}\subset\mathcal{C}^{\prime}$, that is, $\mathcal{E}$ lies in the Müger center of $\mathcal{C}$.
Example 2.4.3.
Assume that $\mathcal{C}$ is a braided finite tensor category, $G$ a finite group and the conditions of 2.3.1 hold, then the sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{G}$
is a braided exact sequence, for $F:\mathcal{C}\to\mathcal{C}_{G}$ the free module functor $F(X)=X\otimes A$ for $A=\text{Fun}(G,\mathbbm{k})$ the algebra object in $\text{Rep}(G)$.
Example 2.4.4.
Assume $\mathcal{C}$ is a braided pointed finite tensor category, and let $G$ be the group of invertible objects of $\mathcal{C}$. Assume $H<G$ is a maximal subgroup such that $\text{Rep}(H)\subset\mathcal{C}^{\prime}$ as a symmetric subcategory. The category $\mathcal{C}_{H}$ is a braided pointed finite tensor category which is non-degenerate (that is, $(\mathcal{C}_{H})^{\prime}=\text{Vec}$), and the group of invertible objects of $\mathcal{C}_{H}$ is $G/H$. Denoting the simple objects of $\mathcal{C}$ by $\{L_{g}\}_{g\in G}$, an invertible object $X\in\mathcal{C}_{H}$ is of the form $\oplus_{g\in B}L_{g}$ (as an object in $\mathcal{C}$) for $B$ a $H$-coset in $G$.
3. The structure of braided simple extensions
Let $\mathcal{C}$ be a braided simple extension of $\mathcal{D}$ with braiding $c$, and assume that $\mathcal{D}$ is not semisimple. Let $Q$ be the unique simple projective object in $\mathcal{C}$ which is not in $\mathcal{D}$, and let $\{L_{i}\}_{i=1}^{n}$ be the (isomorphism classes of) simple objects in $\mathcal{D}$ with projective covers $P_{i}$. The existence of the simple projective object $Q$ and the braiding on $\mathcal{C}$ enables us to understand the structure of $\mathcal{C}$ and $\mathcal{D}$.
3.1. Basic properties of simple extensions
The following lemma is true for a general simple extension $\mathcal{C}$ of $\mathcal{D}$, not necessarily braided:
Lemma 3.1.1.
Let $\mathcal{D}$ be a non semisimple finite tensor category, $\mathcal{C}$ a simple extension of $\mathcal{D}$, and let $Q$ be the simple projective object of $\mathcal{C}$. For a simple object $L\in\mathcal{D}$ with projective cover $P\in\mathcal{D}$ we have $Q\otimes Q:P]=\text{\rm FPdim}_{\mathcal{D}}(L)$.
Proof.
The category $\mathcal{D}$ has a simple extension, hence by definition has a fiber functor $\mathcal{D}\to\text{Vec}$. In particular, $\mathcal{D}$ is integral, and $\text{\rm FPdim}_{\mathcal{D}}(L)=\dim_{\mathbbm{k}}(L)$ (see [EGNO]). $\mathcal{M}=<Q>$ is a $\mathcal{D}$-module category, and $Q$ is projective, hence we have an isomorphism $L\otimes{Q}\cong{\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}Q}$, which gives
$0\neq\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}\text{Hom}_{\mathcal{C}}(Q,Q)={\text{Hom}_{\mathcal{C}}(Q,L\otimes{Q})}=\text{Hom}_{\mathcal{C}}(Q\otimes{Q^{*}},L)=\text{Hom}_{\mathcal{C}}(Q\otimes{Q},L)$
the last equality arises from the fact that $Q$ is self-dual.
Since $L$ is simple, each such nonzero morphism is epi, so we get a morphism $P\to Q\otimes{Q}$, which implies that $P$ is a direct summand of $Q\otimes Q$, because $Q\otimes{Q}$ is projective.
By Schur’s lemma $\dim(\text{Hom}_{\mathcal{C}}(Q,Q))=1$, hence the multiplicity of $P$ in $Q\otimes{Q}$ is exactly $\dim_{\mathbbm{k}}(L)=\text{\rm FPdim}_{\mathcal{D}}(L)$. ∎
By Lemma 3.1.1, we have:
(6) $Q\otimes Q=\oplus_{i=1}^{n}\text{\rm FPdim}(L_{i})P_{i}\oplus rQ$
for some $r\in\mathbb{N}$. By definition, we have $\text{\rm FPdim}(\mathcal{C})=\text{\rm FPdim}(\mathcal{D})+\text{\rm FPdim}(Q)^{2}$, so using Eq. 6 we get that $\text{\rm FPdim}(\mathcal{C})=2\text{\rm FPdim}(\mathcal{D})+r\text{\rm FPdim}(Q)$.
3.2. Degeneracy in braided simple extensions
From here on, let $\mathcal{D}$ be a braided non-semisimple finite tensor category, $\mathcal{C}$ a braided simple extension of $\mathcal{D}$, with a unique simple projective object $Q$. Let $\mathcal{E}_{+}\subset\mathcal{D}$ be the full subcategory generated by objects $X$ which centralize $Q$.
Remark 3.2.1.
Assume we have $X,Y,Z\in\mathcal{C}$ such that $s_{X\otimes Y}=\text{id}_{X\otimes Y}$ and $\psi:X\simeq Z$, then $s_{Z\otimes Y}=\text{id}_{Z\otimes Y}$, by the following diagram:
$X\otimes Y$$X\otimes Y$$Z\otimes Y$$Z\otimes Y$$\scriptstyle{s_{X\otimes Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{s_{Z\otimes Y}}$
which commutes due to the naturality of the braiding $c$.
Proposition 3.2.2.
$\mathcal{E}_{+}$ is a symmetric tensor Serre subcategory of $\mathcal{D}$.
Proof.
For $X,Y\in\mathcal{D}$ centralizing $Q$, we have by Eq. 5 $s_{Q,X\otimes Y}=\text{id}_{Q\otimes X\otimes Y}$, so $X\otimes Y\in\mathcal{E}_{+}$. The object $Q$ is self dual, thus $\mathcal{E}_{+}$ is closed under taking duals (because $s_{Q,X}=\text{id}_{Q\otimes X}$ if and only if $s_{X,Q}=\text{id}_{X\otimes Q}$). This implies that $\mathcal{E}_{+}$ is a rigid monoidal category, and clearly it is abelian and $\mathbbm{k}$-linear as a full subcategory of $\mathcal{D}$.
Consider a short exact sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{D}$. Applying the functor $Q\otimes$- splits this sequence, because $Q$ is projective in $\mathcal{C}$ (and this is a short exact sequence in $\mathcal{C}$ as well), so we have a commutative diagram:
[…]
and the splitting of the two rows imply that $s_{Q,Y}=s_{Q,X}\oplus s_{Q,Z}$, thus $Y$ centralizes $Q$ if and only if $X$ and $Z$ do, which implies that $\mathcal{E}_{+}$ is a Serre subcategory.
Assume $X,Y$ centralize $Q$. By Eq. 4 for the triplet $(Q,X,Y)$ and $s_{Q,Y}=\text{id}_{Q\otimes Y}$ we get:
$s_{Q\otimes X,Y}=a^{-1}_{Q,X,Y}\circ(\text{id}_{Q}\otimes s_{X,Y})\circ a_{Q,X,Y},$
while on the other hand, $Q\otimes X$ is a direct sum of copies of $Q$, so by Remark 3.2.1 $s_{Q\otimes X,Y}=\text{id}_{Q\otimes X\otimes Y}$, which implies that $\text{id}_{Q}\otimes s_{X,Y}=\text{id}_{Q\otimes X\otimes Y}$, thus $s_{X,Y}=\text{id}_{X\otimes Y}$. ∎
By Corollary 8.10.8 in [EGNO], the distinguished invertible object of $\mathcal{C}$ lies in $\mathcal{E}_{+}$.
Corollary 3.2.3.
If $Q\in\mathcal{D}^{\prime}$ then $\mathcal{D}=\mathcal{E}_{+}$ is symmetric.
Proof.
By definition, if $Q\in\mathcal{D}^{\prime}$ then $\mathcal{E}_{+}=\mathcal{D}$, and by 3.2.2, $\mathcal{D}$ is symmetric. ∎
Proposition 3.2.4.
Let $\mathcal{D}$ be a non-semisimple braided finite tensor category, $\mathcal{C}$ a braided simple extension of $\mathcal{D}$ with a simple projective object $Q$, then $s_{Q,Q}\neq\text{id}_{Q\otimes Q}$
Proof.
Assume $s_{Q,Q}=\text{id}_{Q\otimes Q}$, then by Eq. 4 for $X=Y=Z=Q$ we get $s_{Q\otimes Q,Q}=\text{id}_{Q\otimes Q\otimes Q}$. By Eq. 6, we get $s_{Q,P_{i}}=\text{id}_{Q\otimes P_{i}}$, thus $P_{i}\in\mathcal{E}_{+}$ for each $i$. By 3.2.2 $\mathcal{E}_{+}$ is a Serre subcategory, which implies that every indecomposable object $X\in\mathcal{D}$ is also in $\mathcal{E}_{+}$, thus $\mathcal{E}_{+}=\mathcal{D}$.
By 3.2.3, $\mathcal{D}$ is symmetric, and $Q$ centralizes $\mathcal{C}$ (as it centralizes $\mathcal{D}$ and $s_{Q,Q}=\text{id}_{Q\otimes Q}$), thus $\mathcal{C}$ is symmetric.
The category $\mathcal{C}$ is a symmetric finite tensor category over an algebraically closed field of characteristic 0, thus by [D], $\mathcal{C}$ is equivalent to a category of representations of some finite supergroup, but such a category has no simple projective object, unless it is semisimple itself, thus contradicting our assumption. ∎
Theorem 3.2.5.
Let $\mathcal{D}$ be a braided finite tensor category. If $\mathcal{D}$ has a braided simple extension $\mathcal{C}$, then $\mathcal{D}$ is degenerate.
Proof.
Denote $\mathcal{C}=\mathcal{D}\oplus\mathcal{M}$ and let $Q$ be the simple projective object in $\mathcal{M}$. Set $\text{\rm FPdim}(\mathcal{D})=n$ and $\text{\rm FPdim}(Q)=q$ ($q\in\mathbbm{k}^{*}$). By Eq. 6, $q^{2}=n+rq$, thus $\text{\rm FPdim}(\mathcal{C})=2n+rq$. By Eq. 3 and 3.2.2 we get in $\mathcal{C}$:
(7) $n\text{\rm FPdim}(\mathcal{D}^{\prime})=(2n+rq)\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$
which implies that $\text{\rm FPdim}(\mathcal{D}^{\prime})>\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$. If $Q\notin\mathcal{D}^{\prime}$, then $\mathcal{D}^{\prime}$ (the centralizer of $\mathcal{D}$ in $\mathcal{C}$) is a full subcategory of $(\mathcal{D})^{\prime}_{\mathcal{D}}$ (the centralizer of $\mathcal{D}$ in $\mathcal{D}$), with $\text{\rm FPdim}(\mathcal{D}^{\prime})>1$, thus $\text{Vec}\subsetneq(\mathcal{D})^{\prime}_{\mathcal{D}}$, which implies that $\mathcal{D}$ is degenerate. On the other hand, if $Q\in\mathcal{D}^{\prime}$, then by 3.2.3, $\mathcal{D}$ is symmetric, hence degenerate. ∎
Proposition 3.2.6.
$\mathcal{C}^{\prime}=\mathcal{E}_{+}$. In particular, $\mathcal{E}_{+}$ is central in $\mathcal{C}$.
Proof.
We have $\mathcal{E}_{+}\subset\mathcal{D}\subset\mathcal{C}$. We will apply Eq. 3 for the pairs $(\mathcal{E}_{+},\mathcal{D})$ and $(\mathcal{E}_{+},\mathcal{C})$. We get:
(8) $\text{\rm FPdim}(\mathcal{E}_{+})\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})=\text{\rm FPdim}(\mathcal{D})\text{\rm FPdim}(\mathcal{E}_{+}\cap(\mathcal{D})_{\mathcal{D}}^{\prime}),$
(9) $\text{\rm FPdim}(\mathcal{E}_{+})\text{\rm FPdim}(\mathcal{E}_{+}^{\prime})=\text{\rm FPdim}(\mathcal{C})\text{\rm FPdim}(\mathcal{E}_{+}\cap\mathcal{C}^{\prime}).$
On one hand, we have $\mathcal{C}^{\prime}\cap\mathcal{E}_{+}\subset(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$, while on the other hand, for $X\in(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$ we have $s_{X,Y}=\text{id}_{X\otimes Y}$ for all $Y\in\mathcal{D}$ (because $X\in\mathcal{D}^{\prime}$) and for $Y=Q$ (because $X\in\mathcal{E}_{+}$), thus $X\in\mathcal{C}^{\prime}\cap\mathcal{E}_{+}$, hence $\mathcal{C}^{\prime}\cap\mathcal{E}_{+}=(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$, so $\text{\rm FPdim}(\mathcal{E}_{+}\cap(\mathcal{D})_{\mathcal{D}}^{\prime})=\text{\rm FPdim}(\mathcal{E}_{+}\cap\mathcal{C}^{\prime})$. This gives the equality:
(10) $\frac{\text{\rm FPdim}(\mathcal{D})}{\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})}=\frac{\text{\rm FPdim}(\mathcal{C})}{\text{\rm FPdim}(\mathcal{E}_{+}^{\prime})}$
Let us denote $\text{\rm FPdim}(\mathcal{D})=n$, $\text{\rm FPdim}(Q)=q$, $\text{\rm FPdim}(\mathcal{E}^{\prime}_{+})=e_{c}$, $\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})=e_{d}$. By Eq. 6 we have $\text{\rm FPdim}(\mathcal{C})=2n+rq$. In addition, by definition $Q\in\mathcal{E}^{\prime}_{+}$, thus by calculation of Frobenius-Perron dimension we get $e_{c}=e_{d}+q^{2}=e_{d}+n+rq$. Substituting these in Eq. 10 gives:
$\frac{n}{e_{d}}=\frac{2n+rq}{e_{d}+n+rq},$
which translates to the quadratic equation:
$n^{2}+n(rq-e_{d})-e_{d}rq=0.$
This quadratic equation has two solutions: $n=e_{d}$ or n=rqn=-rq. The category $\mathcal{D}$ is integral (it has a fiber functor), and $r,q$ are non-negative [^2] Notice that $q$ is not necessarily integer, but in that case as well n=rqn=-rq is not a valid solution, because $n$ and $r$ are integers., so we must have $n=e_{d}$. This implies that:
$\text{\rm FPdim}(\mathcal{E}^{\prime}_{+})=e_{d}+n+rq=2n+rq=\text{\rm FPdim}(\mathcal{C}),$
thus $\mathcal{E}^{\prime}_{+}=\mathcal{C}$, which in turn implies that $\mathcal{E}_{+}\subseteq\mathcal{C}^{\prime}$. On the other hand, by 3.2.4, $Q\notin\mathcal{C}^{\prime}$, thus $\mathcal{C}^{\prime}\subseteq\mathcal{D}$. This implies that for any $X\in\mathcal{C}^{\prime}$, $X\in\mathcal{D}$ and $s_{Q,X}=\text{id}_{Q\otimes X}$, but this means by definition that $X\in\mathcal{E}_{+}$, thus $\mathcal{C}^{\prime}\subseteq\mathcal{E}_{+}$, completing the proof. ∎
Remark 3.2.7.
Notice that in general, we don’t have $\mathcal{E}_{+}=\mathcal{D}^{\prime}$. For instance, when $\mathcal{D}$ is pointed and $\mathcal{C}$ is non-degenerate, we will prove that $\mathcal{D}$ is slightly degenerate, thus $\mathcal{E}_{+}=\text{Vec}$ while $\mathcal{D}^{\prime}=\text{sVec}$.
4. Braided simple extensions of pointed tensor categories
In this section, we assume that $\mathcal{D}$ is a pointed braided finite tensor category, and $\mathcal{C}$ is a braided simple extension of $\mathcal{D}$. Concrete examples of such categories can be found in Appendix A in [GS]. Denote the simple projective object in $\mathcal{C}$ by $Q$. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a pointed braided fusion subcategory, which we denote by $\mathcal{D}_{f}$. This category corresponds to a pair $(G,q)$ for $G$ a finite abelian group and $q$ a quadratic form on $G$. The group $G$ is simply $\text{Pic}(\mathcal{D})$ (the Picard group), and the squared braiding $s_{L_{g},L_{h}}:L_{g}\otimes L_{h}\to L_{h}\otimes L_{g}$ of two simple objects $L_{g},L_{h}\in\mathcal{D}_{f}$ is multiplication by $b(g,h)=\frac{q(gh)}{q(g)q(h)}$.
From here on, we denote the simple objects of $\mathcal{D}$ by $L_{g}$ ($g\in G$). We denote the projective cover of $L_{g}$ by $P_{g}$, and set $\bar{P}:=\oplus_{g\in\text{Pic}(\mathcal{D})}$. Clearly we have $\bar{P}\otimes L_{g}\cong\bar{P}$ for all $g\in G$.
4.1. The $Q$-braiding group of $\mathcal{C}$
Let $m_{g}$ be the order of $L_{g}$ in $\text{Pic}(\mathcal{D})$ (that is, $m_{g}$ is the minimal integer such that $L_{g}^{\otimes m_{g}}=\text{id}_{\mathcal{D}}$). By the definition of a simple extension, we have
$s_{Q,L_{g}}\in\text{Hom}_{\mathcal{C}}(Q\otimes L_{g},Q\otimes L_{g})=\text{Hom}_{\mathcal{C}}(Q,Q).$
Since $\mathbbm{k}$ is algebraically closed, we get by Schur’s lemma that $s_{Q,L_{g}}$ is simply multiplication by a scalar, which we denote by $\lambda_{g}$.
Applying Eq. 5 to the triplet $(Q,L_{g},L_{h})$, and using the fact that $s_{Q,L_{h}}$ are scalars, we get that $\lambda_{g}\lambda_{h}=\lambda_{gh}$, and $s_{Q,\mathbbm{1}}=1$. In other words, we have a group homomorphism $\psi:\text{Pic}(\mathcal{D})\to\mathbbm{k}^{*}$ given by $\psi(L_{g})=\lambda_{g}$. We define $G_{Q}:=Im(\psi)$ and call this the $Q$-braiding group of $\mathcal{D}$.
Since every $L_{g}$ has a finite order $m_{g}$, by repeatedly using Eq. 5 we see that $\lambda_{g}$ must be a root of unity, of order dividing $m_{g}$, thus $G_{Q}$ is a group of roots of unity of order dividing $|\text{Pic}(\mathcal{D})|$.
What is the kernel of the morphism $\psi$? It contains exactly all the simple objects $L_{g}\in\mathcal{D}$ such that $\lambda_{g}=1$, that is all the simple objects $L_{g}\in\mathcal{D}$ for which $s_{Q,L_{g}}=\text{id}_{Q\otimes L_{g}}$. Each such simple object lies, by definition, in $\mathcal{E}_{+}$, thus $\ker\psi=\text{Pic}(\mathcal{E}_{+})$.
In a way, the group $G_{Q}$ ”encodes” the additional braiding data of the simple extension $\mathcal{C}$ of $\mathcal{D}$.
4.2. Non-semisimple braided near-group categories
Definition 4.2.1.
Let $\mathcal{D}$ be a pointed non-semisimple finite tensor category admitting a fiber functor, and let $G=\text{Pic}(\mathcal{D})$.
  1. (1)
    A simple extension $\mathcal{C}$ of $\mathcal{D}$ is called a non-semisimple near-group category.
  2. (2)
    Denoting the unique simple projective object of $\mathcal{C}$ by $Q$, we say that $\mathcal{C}$ has a generalized fusion rule $(G,r)$ for $r$ such that Eq. 6 holds.
  3. (3)
    If $\mathcal{D}$ and $\mathcal{C}$ are braided, and the restriction of the braiding of $\mathcal{C}$ to $\mathcal{D}$ coincide with the braiding of $\mathcal{D}$, we say that $\mathcal{C}$ is a braided non-semisimple near-group category.
Proposition 4.2.2.
For a braided non-semisimple near-group category $\mathcal{C}$ with a generalized fusion rule $(G,r)$, if $r>0$ then $\mathcal{E}_{+}=\mathcal{D}$, i.e. $\mathcal{D}$ is symmetric and $\mathcal{D}^{\prime}=\mathcal{C}$.
Proof.
Consider the morphism $s_{Q\otimes Q,L_{g}}$. On one hand, by Eq. 4 for the triplet $(Q,Q,L_{g})$ we get (see also Remark 3.2.1):
$s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}=\lambda_{g}^{2}\text{id}_{\bar{P}}\oplus(\oplus_{r}\lambda_{g}^{2}\cdot\text{id}_{Q})$
We use here the isomorphism $Q\otimes L_{g}\cong Q$. On the other hand, using the naturality of the square braiding we have:
$s_{Q\otimes Q,L_{g}}=s_{\bar{P},L_{g}}\oplus(\oplus_{r}s_{Q,L_{g}})=s_{\bar{P},L_{g}}\oplus(\oplus_{r}\lambda_{g}\text{id}_{Q})$
Comparing these two expressions and using our assumption that $r>0$, we conclude that $\lambda_{g}=1$ for all $g$. This implies that $L_{g}\in\mathcal{E}_{+}$ for all $g\in G$. By 3.2.2, $\mathcal{E}_{+}$ is a Serre subcategory of $\mathcal{D}$, thus $\mathcal{E}_{+}=\mathcal{D}$, hence $\mathcal{D}$ is symmetric. By 3.2.6 we get that $\mathcal{D}^{\prime}=\mathcal{C}$. ∎
Theorem 1. Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$. In particular, a non-semisimple braided near-group category is weakly integral.
Proof of 1.
Let $\mathcal{C}$ be a braided non-semisimple near-group category with a generalized fusion rule $(G,r)$, and assume that $r>0$. By 4.2.2, we get that $\mathcal{D}$ is a symmetric subcategory of $\mathcal{C}$ and $Q\in\mathcal{D}^{\prime}$. By [D], since $\mathcal{D}$ is not semisimple, it is braid equivalent to $\textrm{sRep}(G\ltimes W,u)$, a category of finite supermodules of the supergroup $G\ltimes W$, with $u\in G$ a central element of order $2$ acting as the parity. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a fusion subcategory, which is braid equivalent to $\text{Vec}_{G\times\mathbb{Z}/2\mathbb{Z}}$ (the central element $u$ generates $\mathbb{Z}/2\mathbb{Z}$). In particular, there exist a simple object $S\in\mathcal{D}_{f}\subset\mathcal{D}$ such that $S\otimes S=\mathbbm{1}$ and $c_{S,S}=-\tau_{S,S}$ for $\tau_{X,Y}$ the flip morphism $\tau_{X,Y}:X\otimes Y\to Y\otimes X$.
The objects $S$ and $Q$ are simple objects, and $S$ is invertible, thus we have an isomorphism $\psi:Q\otimes S\simeq Q$, and the map $c_{S,Q}:S\otimes Q\to Q\otimes S$ is given by some scalar $\lambda$, that is $c_{S,Q}=\lambda\tau_{S,Q}$. The tensor product in $\mathcal{C}$ is compatible with the $\mathcal{D}$-bimodule structure of $\mathcal{M}$, which in turn is compatible with the $\mathcal{D}$-bimodule structure of Vec, thus for $a$ the associativity constraint in $\mathcal{C}$ we have $a_{S,S,Q}=\text{id}_{S\otimes S\otimes Q}$, $a_{S,Q,S}=\text{id}_{S\otimes Q\otimes S}$ and $a_{Q,S,S}=\text{id}_{Q\otimes S\otimes S}$.
Consider the following diagram:
$S\otimes Q\otimes S$$S\otimes S\otimes Q$$Q\otimes S\otimes S$$\mathbbm{1}\otimes Q$$Q\otimes\mathbbm{1}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes c_{S,Q}}$$\scriptstyle{c_{S\otimes S,Q}}$$\scriptstyle{c_{\mathbbm{1},Q}}$
the upper triangle is commutative by the axiom of the braiding (suppressing the associativity constraints which are identities as explained above), and the lower square commutes by the naturality of the braiding. This diagram implies that $\lambda^{2}=1$, thus in particular $\lambda^{-1}=\lambda$.
Now consider the following diagram:
$S\otimes Q\otimes S$$Q\otimes S\otimes S$$Q\otimes S\otimes S$$S\otimes Q$$Q\otimes S$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes\psi}$$\scriptstyle{\text{id}_{Q}\otimes c_{S,S}}$$\scriptstyle{\psi\otimes\text{id}_{S}}$$\scriptstyle{c_{S,Q}}$
this diagram commutes due to the naturality of the braiding, so we get an equality:
$\lambda(\psi\otimes\text{id}_{S})\circ(\text{id}_{Q}\circ\tau_{S,S})\circ(\tau_{S,Q}\otimes\text{id}_{S})=(\lambda\tau_{S,Q})\circ(\text{id}_{S}\otimes\psi)$
which must imply that $\lambda=-\lambda$, giving a contradiction, thus $r=0$.
By Eq. 6, if $r=0$ then $Q\otimes Q=\oplus_{g\in G}P_{g}$ for $P_{g}$ the projective cover of the simple object $L_{g}$, and all the projective covers of simple objects have the same Frobenius-Perron dimension, thus $\text{\rm FPdim}(\mathcal{C})=2|G|\text{\rm FPdim}(P_{e})$ for $P_{e}$ the projective cover of $\mathbbm{1}$, hence $\mathcal{C}$ is weakly integral. ∎
As a result of 1, when we consider non-semisimple braided near-group categories $\mathcal{C}$, the generalized fusion rule is always $(\text{Pic}(\mathcal{D}),0)$, so as an abbreviation, we will simply say that $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category.
4.3. De-equivariantization of near-group categories
In this subsection, we prove that every $\text{Pic}(\mathcal{D})$-braided near-group category is ”an extension” of a non-degenerate braided non-semisimple near-group category by some symmetric fusion category, that is, there is a braided exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{nd}$
for $\mathcal{C}_{nd}$ a non-degenerate braided non-semisimple near-group category. This implies that the classification of braided non-semisimple near-group categories restricts to the classification of non-degenerate braided non-semisimple near-group categories, in a way.
Proposition 4.3.1.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category for some $\mathcal{D}$. Assume that for $H<\text{Pic}(\mathcal{D})$ we have $\text{Rep}(H)\subset\mathcal{C}^{\prime}$, then $\mathcal{C}_{H}$ is a braided non-semisimple near-group category.
Proof.
By 2.3.1 $\mathcal{C}_{H}$ is a well-defined braided finite tensor category, and the free module functor $F:\mathcal{C}\to\mathcal{C}_{H}$ is a braided tensor functor. This functor is surjective and normal by [BN3].
Let $Y\in\mathcal{C}_{H}$ be a simple object. By Lemma 2.1 in [BN2], $Y$ is a quotient of $F(X)$ for some $X\in\mathcal{C}$, that is we have a surjective morphism $\psi:F(X)\to Y$. Since $Y$ is simple and $F$ is a tensor functor (hence additive), we may assume that $X\in\mathcal{C}$ is indecomposable, and we can choose $X$ such that $\psi:F(X)\to L$ is non-zero and $X$ has minimal Loewy length. Let $L$ be the top of $X$, then $L=Q$ or $L\in\mathcal{D}$, and we have a surjective morphism $\varphi:X\to L$.
Assume $L\in\mathcal{D}$. We have an exact sequence in $\mathcal{D}$:
$0\to Z\to X\xrightarrow{\varphi}L\to 0$
and $\text{\rm FPdim}(L)=1$, thus by 2.3.3, $F(L)$ is also invertible, and $F(\varphi):F(X)\to F(L)$ is a surjective morphism (because $F$ is exact). If $F(L)\simeq Y$ we are done, otherwise $\varphi|_{F(Z)}:F(Z)\to Y$ must be non-zero, but the Loewy length of $Z$ is smaller than that of $X$, which contradicts the choice of $X$.
If $L=Q$, by the assumption that $X$ is indecomposable, we must have $X=Q$. By [EO], surjective tensor functors map projective objects to projective objects, thus $F(X)=Q\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$ for $A:=\text{Fun}(H,\mathbbm{k})$ the algebra object in $\text{Rep}(H)$. By Lemma 7.8.12 in [EGNO], $\text{Hom}_{A}(Q\otimes A,Y)=\text{Hom}_{\mathcal{C}}(Q,\bar{Y})$ for $\bar{Y}$ the $A$-module $Y$ considered as an object in $\mathcal{C}$, thus as an object in $\mathcal{C}$, we have $Y=Z\oplus Q^{\oplus m}$ for some $m\in\mathbb{N}$, $Z\in\mathcal{D}$.
The category $\textrm{Mod}_{\mathcal{C}}(A)$ is additive, and as an object in $\mathcal{C}$, $A$ is a direct sum of simple objects in $\mathcal{D}$, thus $Q\otimes A$ is a direct sum of copies of $Q$, while $Z\otimes A\in\mathcal{D}$. Since we assume that $Y$ is simple as an $A$-module, this implies that $Z=0$.
The algebra $A$ is separable (as an algebra object in $\text{Rep}(H)$), so $Y$ is a direct summand of $Y\otimes A$ in $\textrm{Mod}_{\mathcal{C}}(A)$, so to prove that $Y$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, it is enough to prove that $Y\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, but this is clear because of Lemma 7.8.12 in [EGNO] and the fact that $\bar{Y}$ is a direct sum of copies of $Q$. In addition, it is clear from the proof that such $Y$ is uniquely determined by the fact that $L=Q$.
We conclude that $\textrm{Mod}_{\mathcal{C}}(A)$ is a braided, with a unique simple projective object, and all the other objects are invertible, hence it is a braided non-semisimple near-group category. ∎
Corollary 4.3.2.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group, $H<\text{Pic}(\mathcal{D})$ such that $\text{Rep}(H)\subset\mathcal{C}^{\prime}$, and for $X\in\mathcal{C}_{H}$ denote by $\bar{X}$ the underlying object in $\mathcal{C}$.
  1. (1)
    If $Y\in\mathcal{C}_{H}$ is the unique simple projective object, then $\bar{Y}=Q^{\oplus m}$ for $m\in\mathbb{N}$.
  2. (2)
    If $Y\in\mathcal{C}_{H}$ is invertible, then $\bar{Y}$ is a direct sum of invertible objects in $\mathcal{D}$.
Proof.
The first claim follows directly from the proof of 4.3.1. The second claim follows from Example 2.4.4. ∎
Proposition 4.3.3.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category, then there exist $H<\text{Pic}(\mathcal{D})$ such that $\mathcal{C}_{H}$ is a non-degenerate braided near-group category.
Proof.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, $\mathcal{E}_{+}$ is a central symmetric subcategory of $\mathcal{C}$ and $\mathcal{D}$, thus in particular it is a symmetric pointed tensor category. Set $H:=\text{Pic}(\mathcal{E}_{+})$, then $\text{Rep}(H)$ is also symmetric and central in $\mathcal{D}$ and $\mathcal{C}$. By 4.3.1, $\mathcal{C}_{H}$ is a braided non-semisimple near-group category, so we need only to prove that $\mathcal{C}_{H}$ is non-degenerate. Set $A:=\text{Fun}(H,\mathbbm{k})$.
Let $Y\in\mathcal{C}_{H}$ be the unique simple projective object, and let $L\in\mathcal{C}_{H}$ be a simple invertible object. By 4.3.2, we have $\bar{Y}=Q^{\oplus m}$ for some $m\in\mathbb{N}$ and $\bar{L}=\oplus_{f\in B}L_{f}$ for $B$ a coset of $H$ in $G$. Assume that $\tilde{s}_{Y,L}=\text{id}_{Y\otimes_{A}L}$ for $\tilde{s}$ the squared braiding in $\mathcal{C}_{H}$. By 2.3.1 we have a commutative diagram:
$\bar{Y}\otimes\bar{L}$$\bar{L}\otimes\bar{Y}$$Y\otimes_{A}L$$L\otimes_{A}Y$$\scriptstyle{s_{\bar{Y},\bar{L}}}$$\scriptstyle{\tilde{s}_{Y,L}}$
thus $s_{\bar{Y},\bar{L}}=\text{id}_{\bar{Y}\otimes\bar{L}}$. By the definition of $\mathcal{E}_{+}$, since $\bar{Y}=Q^{\oplus m}$, we have $L_{f}\in\mathcal{E}_{+}$ for all $f\in B$. This implies (by Example 2.4.4) that $B$ is the trivial coset, that is $\bar{L}=\oplus_{h\in H}L_{h}$, hence $L=\mathbbm{1}_{\mathcal{C}_{H}}$, so in $\mathcal{C}_{H}$ the object $Y$ is centralized only by $\mathbbm{1}_{\mathcal{C}_{H}}$, thus $\mathcal{C}_{H}$ is non-degenerate as required. ∎
Theorem 2. Let $\mathcal{C}$ be a braided non-semisimple near-group category, then there exist a non-degenerate braided non-semisimple near-group category $\mathcal{D}$ and a finite group $G$ such that $\mathcal{C}$ fits in a modularization exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}.$
Proof of 2.
Let $\mathcal{C}$ be a braided non-semisimple near-group category, i.e. a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$. By 1, $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, the category $\mathcal{E}_{+}$ generated by objects in $\mathcal{D}$ that cetralize the unique simple projective object $Q\in\mathcal{C}$ is a symmetric subcategory of $\mathcal{D}$ which is central in $\mathcal{C}$. Hence, by 4.3.3, taking $H:=\text{Pic}(\mathcal{E}_{+})$ we get that $\mathcal{C}_{H}$ is a non-degenerate $G_{Q}$-braided near-group category and we get a modularization exact sequence:
$\text{Rep}(\text{Pic}(\mathcal{E}_{+}))\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{H}.$
4.4. Non-degenerate near-group categories
As a result of 2, in order to fully classify braided non-semisimple near-group categories, we need to classify non-degenerate braided non-semisimple near-group categories. This section is devoted to this classification.
Proposition 4.4.1.
If $\mathcal{C}$ is a non-degenerate braided non-semisimple near-group category with a generalized fusion rule $(G,0)$, then $G=\mathbb{Z}_{2}$.
Proof.
By 3.2.2, since $\mathcal{C}$ is non-degenerate we have $\text{\rm FPdim}(\mathcal{E}_{+})=1$, thus by Eq. 7 (together with $r=0$) we get that $\text{\rm FPdim}(\mathcal{D}^{\prime})=2$, that is $\mathcal{D}$ is slightly degenerate.
Let $L_{g}\in\mathcal{D}$ be a non-trivial simple object, and denote $s_{Q,L_{g}}=\lambda_{g}$, then . By Eq. 4 for the triplet $(Q,Q,L_{g})$ we get that $s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}$. On the other hand, by Eq. 6, the naturality of the braiding, and the fact that $s_{Q\otimes Q,L_{g}}$ is a scalar map, we get that $s_{\bar{P},L_{g}}=\lambda_{g}^{2}\bar{P}\otimes L_{g}$. In particular, for $P_{e}$ the projective cover of $\mathbbm{1}$ we have $s_{P_{e},L_{g}}=\lambda_{g}^{2}\text{id}_{P_{e}\otimes L_{g}}$. $\mathcal{C}$ is non-degenerate, thus unimodular (by [EGNO]), so $\mathbbm{1}$ is the socle of $P_{e}$, thus $\lambda^{2}_{g}\text{id}_{L_{g}}=s_{\mathbbm{1},L_{g}}=\text{id}_{L_{g}}$, hence $\lambda_{g}^{2}=1$.
We conclude that for each projective cover $P_{h}$ of $L_{h}\in\text{Pic}(\mathcal{D})$ we have $s_{P_{h},L_{g}}=\text{id}_{P_{h}\otimes L_{g}}$, thus for each projective object $P\in\mathcal{D}$ we must have $s_{P,L_{g}}=\text{id}_{P\otimes L_{g}}$. Let $X\in\mathcal{D}$ be indecomposable, then $X$ has a projective cover $P^{\prime}$, a surjective map $\pi:P^{\prime}\to X$, and we get a commutative diagram:
$P^{\prime}\otimes L_{g}$$P^{\prime}\otimes L_{g}$$X\otimes L_{g}$$X\otimes L_{g}$$\scriptstyle{s_{P^{\prime}\otimes L_{g}}}$$\scriptstyle{\pi\otimes\text{id}_{L_{g}}}$$\scriptstyle{\pi\otimes\text{id}_{L_{g}}}$$\scriptstyle{s_{X,L_{g}}}$
which implies that $s_{X,L_{g}}=\text{id}_{X\otimes L_{g}}$ as well. Overall, we get that for $L_{g}$ a non-trivial simple object in $\mathcal{D}$, $L_{g}\in\mathcal{D}^{\prime}$, but $\mathcal{D}^{\prime}$ is slightly degenerate, thus $L_{g}$ must be the unique non-trivial invertible element in $\mathcal{D}^{\prime}\simeq\text{sVec}$, hence $G=\mathbb{Z}_{2}$ as required.
Theorem 3. Every non-degenerate non-semisimple braided near-group category $C$ is a braided simple extension of a category $\mathcal{D}$ braid equivalent to $\textrm{sRep}(W\oplus W^{*})$, with $\textrm{sRep}(W)$ a Lagrangian subcategory, for some purely odd supervector space $W$.
Proof of 3.
Let $\mathcal{C}$ be a non-degenerate, braided non-semisimple near-group category with fusion rule $(G,0)$. By 4.4.1, we have $G=\mathbb{Z}_{2}$. This implies that $\mathcal{D}$ is a braided finite tensor category with only two simple objects, both invertible. Since $\mathcal{D}_{f}=\text{sVec}$ is a symmetric subcategory of $\mathcal{D}$, there is some Lagrangian subcategory $\mathcal{D}_{0}$ in $\mathcal{D}$, and we must have $\text{sVec}\subset\mathcal{D}_{0}$, so by [D], $\mathcal{D}_{0}$ is equivalent to a category of representations of some supergroup, and since $|G|=2$ this supergroup must be $\wedge W$ for some purely odd supervector space $W$, that is $\mathcal{D}_{0}\cong\textrm{sRep}(W)$. Now using [GS] we get the required result. ∎
In 1 we proved that every braided non-semisimple near-group category is weakly integral. Using 2 and 3 we get a stronger result:
Corollary 4.4.2.
A braided non-semisimple near-group category $\mathcal{C}$ is non-integral weakly integral.
Proof.
by 1 we know that $\mathcal{C}$ is weakly integral, so we need to prove it is not integral. By 2, we have an exact sequence:
$0\to\text{Rep}(G)\to\mathcal{C}\to\bar{\mathcal{C}}\to 0$
for $\bar{\mathcal{C}}$ a non-degenerate braided non-semisimple near-group category. By 3, $\bar{\mathcal{C}}$ is a simple extension of $\bar{\mathcal{D}}$ which is braid equivalent to $\textrm{sRep}(W\oplus W^{*})$. A direct calculation shows that $\text{\rm FPdim}(\bar{\mathcal{D}})=2^{2\dim(W)+1}$, so denoting by $Q_{0}$ the unique simple projective object of $\bar{\mathcal{C}}$, we have $\text{\rm FPdim}(Q_{0})=\sqrt{2}\cdot 2^{\dim(W)}$. Denoting by $Q$ the unique simple object in $\mathcal{C}$, we know that the functor $\mathcal{C}\to\bar{\mathcal{C}}$ is the free module functor and it maps $Q$ to $Q_{0}$. Since this is a surjective functor, by [EO] it preserve Frobenius-Perron dimensions, thus $\text{\rm FPdim}(Q)=\sqrt{2}\cdot^{\dim(W)}\notin\mathbb{N}$. ∎
Using 2 and 3 we get the following result:
Proposition 4.4.3.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided non-semisimple near-group category, then $\mathcal{D}$ has a simple object $S$ such that $S\otimes S=\mathbbm{1}$ and $S\in\mathcal{D}^{\prime}$.
Proof.
By 2 and 3, for $\mathcal{E}_{+}=\mathcal{C}^{\prime}$ we have a modularization exact sequence
$\text{Rep}(\text{Pic}(\mathcal{E}_{+}))\to\mathcal{C}\xrightarrow{F}\mathcal{C}_{\text{Pic}(\mathcal{E}_{+})}$
and $\mathcal{C}_{\text{Pic}(\mathcal{E}_{+})}$ is a non-degenerate braided non-semisimple near-group category. Denote $G:=\text{Pic}(\mathcal{C})$, $H:=\text{Pic}(\mathcal{E}_{+})$ (so $H<G)$, and denote the simple invertible objects of $\mathcal{C}$ by $L_{g}$ for $g\in G$.
$\mathcal{C}_{H}$ has a unique non-trivial invertible object $\tilde{S}$. The functor $F$ is dominant, so we have $X\in\mathcal{C}$ such that $\tilde{S}$ is a subobject of $F(X)$. Let $L_{g}$ be the socle of $F(X)$, then $F(L_{g})\simeq\tilde{S}$, hence $g\notin H$.
The functor $F$ is a tensor functor, so $F(L_{g}\otimes L_{g})\simeq F(L_{g})\otimes F(L_{g})\simeq\tilde{S}\otimes\tilde{S}=\mathbbm{1}$, thus $g^{2}\in H$.
$G$ is an abelian group, so $(gH)^{2}=g^{2}H=H$, that is, $H$ is a subgroup of index $2$, thus there is $g_{0}\in G\setminus H$ such that $g_{0}^{2}=1_{G}$. Set $S:=L_{g_{0}}$, then $S\otimes S=\mathbbm{1}_{\mathcal{C}}$.
By 3, the category $\mathcal{C}_{H}$ has a braided subcategory $\tilde{\mathcal{D}}$ which is tensor equivalent to $\textrm{sRep}(W\oplus W^{*})$ for some purely odd supervector space $W$, and by [GS], $\tilde{S}\in\tilde{\mathcal{D}}^{\prime}$.
By the proof of 4.3.1, the functor $F:\mathcal{C}\to\mathcal{C}_{H}$ restricts (as a braided tensor functor) to $F_{\mathcal{D}}:\mathcal{D}\to\tilde{\mathcal{D}}$, thus by 2.3.1, we have $s_{X,S}=\text{id}_{X\otimes S}$ for all $X\in\text{Pic}(\mathcal{D})$. ∎
This proposition, along with the previous results, gives the following:
Theorem 5. Every braided non-semisimple near-group category $\mathcal{C}$ corresponds to a pair $(G,n_{p})$ such that:
  1. (1)
    $G$ is a group of order $2^{n_{g}}$ with a central element of order $2$.
  2. (2)
    $n_{p}\in\mathbb{N}$ and for $P_{\mathbbm{1}}$ the projective cover of $\mathbbm{1}_{\mathcal{C}}$, $\dim(P_{e})=2^{n_{p}}$ as a vector space.
  3. (3)
    $n_{p}+n_{g}$ is an odd number.
Proof.
By definition, $\mathcal{C}$ is a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$, and denote $G:=\text{Pic}(\mathcal{D})$. By 4.4.3, $G$ has a central element of order $2$. By the proof of 4.4.2, there exist $l\in\mathbb{N}$ such that $\text{\rm FPdim}(Q)=\sqrt{2}\cdot 2^{l}$. By 1 and Eq. 6 we get that $\text{\rm FPdim}(\mathcal{D})=2^{2l+1}$. On the other hand, $\mathcal{D}$ is pointed, so denoting its simple objects by $L_{g}$ and their projective covers by $P_{g}$, for $g\in G$, we know that $\text{\rm FPdim}(P_{g})$ is invariant to $g$, and $\text{\rm FPdim}(\mathcal{D})=|G|\cdot\text{\rm FPdim}(P_{e})$ (for $e\in G$ the unit). This imply that there exist $n_{g}\in\mathbb{N}$ such that $|G|=2^{n_{g}}$ and $\text{\rm FPdim}(P_{e})=2^{2l+1-n_{g}}$, that is $n_{p}=2l+1-n_{g}$. Since $\mathcal{D}$ is integral, $\dim(P_{e})=\text{\rm FPdim}(P_{e})=2^{n_{p}}$, and $n_{p}+n_{g}=2l+1$ is an odd number. ∎