### v2

# On braided simple extensions and braided non-semisimple near-group categories

Abstract.
We study simple extensions of pointed finite tensor categories, that is, tensor categories $\mathcal{C}$ admitting an abelian decomposition $\mathcal{C}\cong\mathcal{D}\oplus\mathcal{M}$ where $\mathcal{D}$ is a pointed tensor subcategory and $\mathcal{M}$ has a unique simple projective object. Such categories provide a natural generalization of near-group categories. Our results concern the braided case.
We prove that every non-degenerate braided non-semisimple near-group category is a braided simple extension of $\textrm{sRep}(W\oplus W^{*})$ with non-trivial braiding for which $\textrm{sRep}(W)$ is Lagrangian. Moreover, any braided non-semisimple near-group category $\mathcal{C}$ arises canonically as an extension of such a category by $\text{Rep}(G)$, where $G$ is the Picard group of a symmetric subcategory determined by the unique simple projective object of $\mathcal{C}$.
Keywords: finite tensor categories, near-group categories, non-semisimple categories, braided tensor categories
MSC 2020: 18M05, 18M15.
1. Introduction
Let $\mathcal{D}$ be a finite tensor category; that is, an abelian, $\mathbbm{k}$-linear, rigid monoidal category with a finite number of (isomorphism classes of) simple objects, where $\mathbbm{1}$ is a simple object and each object has a projective cover. For a monoidal category $\mathcal{C}$, a category $\mathcal{M}$ is called a left $\mathcal{C}$-module category if there exists a bifunctor $\tilde{\otimes}:\mathcal{C}\times\mathcal{M}\to\mathcal{M}$ and a natural isomorphism
$m_{X,Y,M}:(X\otimes Y)\tilde{\otimes}M\to X\tilde{\otimes}(Y\tilde{\otimes}M)$
for $X,Y\in\mathcal{C}$ and $M\in\mathcal{M}$ which is compatible with the associativity constraint of $\mathcal{C}$ and $M\mapsto\mathbbm{1}\tilde{\otimes}M$ is an autoequivalence of $\mathcal{M}$.
A classical result in the theory of finite tensor categories is that if $\mathcal{D}$ has a fiber functor, then Vec is a semisimple $\mathcal{D}$-bimodule category, and $\mathcal{C}:=\mathcal{D}\oplus\text{Vec}$ is also a $\mathcal{D}$-bimodule category (see [1]). A natural question that arises is: when can we define a monoidal structure on $\mathcal{C}$. To be more precise, we ask:
Question 1.0.1.
For a finite tensor category $\mathcal{D}$ with a fiber functor $F:\mathcal{D}\to\text{Vec}$, is there a category $\mathcal{M}$ which is equivalent to Vec as a $\mathcal{D}$-bimodule category, such that $\mathcal{C}:=\mathcal{D}\oplus\mathcal{M}$ has a structure of a finite tensor category (which is compatible with the $\mathcal{D}$-action on $\mathcal{M}$)?
When the answer to this question is yes, we say that $(\mathcal{C},\mathcal{D},\mathcal{M})$ is a simple extension of $\mathcal{D}$. Usually, we will omit $\mathcal{M}$ and simply say that $\mathcal{C}$ is a simple extension of $\mathcal{D}$. Notice that, by definition, if $\mathcal{C}$ is a simple extension of $\mathcal{D}$, then $\mathcal{C}$ has a simple projective object $Q$, the unique simple object in $\mathcal{M}$.
In fusion categories, there is a family of categories called near-group categories, which were introduced in [2] (though classification of such categories appeared even earlier in [3]). A near-group category is a fusion category $\mathcal{C}$ with a unique non-invertible object $Q$. Denoting the group of simple invertible objects of $\mathcal{C}$ by $G=\{L_{i}\}$, then a key parameter in the structure of $\mathcal{C}$ is the parameter $r\in\mathbb{N}$ such that $Q\otimes Q\cong\oplus_{i}L_{i}\oplus rQ$. This near-group category is usually denoted by $\mathcal{C}(G,r)$, and we say that $\mathcal{C}$ has a fusion rule $(G,r)$.
There are many known results regarding near-group categories. For instance, in their work, [3] proved that every near-group category with fusion rule $(G,0)$ must be non-strict, that is, must have non-trivial associativity constraint (they also determine the structure constants of the possible associativity constraints in such categories). In addition, [4] classified braided near-group categories with fusion rule $(G,0)$, and their braid structures (including the structure of the underlying group $G$), and [5] classified braided near-group categories with fusion rule $(G,r)$ for $r>0$. These results strongly rely on two factors:
  • The semisimplicity of near-group categories, which enables translating the diagrams defining (braided) fusion categories to equations describing the fusion rules $(G,r)$, due to their unique structure (as a fusion category with a unique non-invertible object).
  • The classification of pointed fusion categories, as categories of the form $\text{Vec}^{\omega}_{G}$ of $G$-graded vector spaces, with associativity constraint corresponding to a $3$-cocycle $\omega\in H^{3}(G,\mathbbm{k}^{*})$.
A natural way to generalize near-group categories to the class of finite tensor categories is:
Definition 1.0.2.
A non-semisimple finite tensor category $\mathcal{C}$ is called a non-semisimple near-group category if $\mathcal{C}$ is a simple extension of a non-semisimple pointed finite tensor category $\mathcal{D}$.
Extending the notion from fusion categories, we say that a non-semisimple near-group category $\mathcal{C}$ has generalized fusion rule $(G,r)$ if $G$ is the group of invertible objects of $\mathcal{C}$ (that is, the Picard group of $\mathcal{D}$, of which $\mathcal{C}$ is its simple extension) and $r$ is the number of copies of the unique simple projective object $Q$ in the tensor product $Q\otimes Q$.
[…]
In this paper, we wish to take a first step toward classifying non-semi-simple near-group categories, by focusing on braided non-semisimple near-group categories, i.e. braided simple extensions of braided pointed finite tensor categories, admitting a fiber functor. Such a classification should be easier, due to the classification of braided pointed finite tensor categories with a fiber functor, which is given in [7].
Formally, in this work we consider triplets of simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$ for which $\mathcal{D}$ is braided, $\mathcal{M}$ is equivalent as a $\mathcal{D}$-bimodule category to Vec and $\mathcal{C}$ is the simple extension of $\mathcal{D}$, such that $\mathcal{C}$ is braided and $\mathcal{D}$ is a braided tensor subcategory of $\mathcal{C}$. An example for such a category is given in [8]. On the other hand, there are braided finite tensor categories with simple extensions that are not braided - for instance $\text{Rep}(H_{4})$ is braided but $\text{Rep}(A^{\prime\prime}_{C_{4}})$ is not, because $A^{\prime\prime}_{C_{4}}$ has no quasi-triangular structure (see [6]).
In the fusion case, [5] proves that there are braided near-group categories with fusion rule $(G,r)$ for $r>0$. In this paper, we prove that this fails in non-semisimple near-group categories. That is, we prove the following theorem:
Theorem 1.
Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$, and consequently $\mathcal{C}$ is weakly integral.
In their work ([9], [10]), the authors defined the notion of a central exact sequence of finite tensor categories as an exact sequence $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ of finite tensor categories with exact tensor functors $i:\mathcal{E}\to\mathcal{C}$ and $F:\mathcal{C}\to\mathcal{D}$ such that $i$ is a full embedding, $F$ is surjective and normal, meaning that for every $Y\in\mathcal{C}$ there exists a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the maximal subobject of $F(X)$ that is a direct sum of copies of $\mathbbm{1}_{\mathcal{D}}$, and $F\circ i$ is a fiber functor of $\mathcal{E}$. Such an exact sequence is called braided if the categories $\mathcal{E},\mathcal{C},\mathcal{D}$ and the functors $i,F$ are braided, and it is called a modularization exact sequence if $\mathcal{D}$ is modular (i.e, non-degenerate). In this paper, we prove that braided non-semisimple near-group categories are ”extensions” of non-degenerate braided non-semisimple near-group categories by symmetric fusion categories. Formally, we prove:
Theorem 2.
Let $\mathcal{C}$ be a braided non-semisimple near-group category, then there exist a non-degenerate braided non-semisimple near-group category $\mathcal{D}$ and a finite group $G$ such that $\mathcal{C}$ fits in a modularization exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}.$
The proof of this theorem is constructive, and $(\mathcal{D},G)$ are determined by $\mathcal{C}$. As a result of this theorem, classifying braided non-semisimple near-group categories restricts to classification of non-degenerate braided non-semisimple near-group categories.
In their work, [8] classified non-degenerate braided finite tensor categories with a Lagrangian subcategory braid equivalent to $\textrm{sRep}(W)$ (the category of supermodules of the exterior algebra over a purely odd supervector space $W$), and they proved that some of these categories are simple braided extensions of $\textrm{sRep}(W\oplus W^{*})$ (though they do not use the term ”braided simple extension”). In this paper, we prove that these examples are, in a way, unique. To be exact, we prove:
Theorem 3.
Every non-degenerate non-semisimple braided near-group category $C$ is a braided simple extension of a category $\mathcal{D}$ braid equivalent to $\textrm{sRep}(W\oplus W^{*})$, with $\textrm{sRep}(W)$ a Lagrangian subcategory, for some purely odd supervector space $W$.
Notice that $\mathcal{D}$ is equivalent to $\textrm{sRep}(W\oplus W^{*})$, with a non-symmetric braid structure (as the theorem states, $\textrm{sRep}(W)$ is the maximal symmetric subcategory of $\mathcal{D}$, i.e. the maximal subcategory in which the restriction of the braiding is symmetric).
The combination of 2 and 3 gives a complete description of the class of braided non-semisimple near-group categories. This classification give two results:
Corollary 4.
Every braided non-semisimple near-group category is not integral.
Theorem 5.
Every braided non-semisimple near-group category $\mathcal{C}$ corresponds to a pair $(G,n_{p})$ such that:
  1. (1)
    $G$ is a group of order $2^{n_{g}}$ with a central element of order $2$.
  2. (2)
    $n_{p}\in\mathbb{N}$ and for $P_{\mathbbm{1}}$ the projective cover of $\mathbbm{1}_{\mathcal{C}}$, $\dim(P_{e})=2^{n_{p}}$ as a vector space.
  3. (3)
    $n_{p}+n_{g}$ is an odd number.
This result is a slightly weaker version of the result given in [4] concerning braided (fusion) near-group categories. In particular, this Theorem gives a categorical proof to the result of [6] stated above that there is no braided simple extension to $\text{Rep}(H_{4})$, in $\text{Rep}(H_{4})$ we have $n_{g}+n_{p}=2$.
1.1. Structure of the paper
In Section 2, we give the necessary preliminaries, including results concerning the Müger centralizer in braided finite tensor categories, which will be used extensively.
In Section 3 we consider general braided simple extensions $(\mathcal{C},\mathcal{D},\mathcal{M})$, and develop some basic tools that will be required to the study of braided non-semisimple near-group categories. In Section 4 we focus on braided non-semisimple near-group categories.
Acknowledgments
I’m grateful to Inna Entova-Aizenbud for helpful discussions. This research was supported by the Israel Science Foundation grant ISF No. 1362/23 (PI: Inna Entova-Aizenbud).
2. Preliminaries
In this section, we recall some basic background on finite tensor categories and braiding of tensor categories. Throughout this paper, we work over $\mathbbm{k}$, an algebraically closed field of characteristic zero, and all categories are assumed to be abelian and $\mathbbm{k}$-linear.
2.1. Notation and terminology
By a finite tensor category $\mathcal{C}$, we mean a $\mathbbm{k}$-linear, abelian, monoidal category with finitely many simple objects (up to isomorphism), such that each simple object has a projective cover, and $\mathbbm{1}$ is a simple object. Throughout this paper, we assume all the tensor categories have trivial unit constraints, but not necessarily trivial associativity constraint. $\mathcal{C}$ is called semisimple if every object is a direct sum of simple objects.
For an abelian category $\mathcal{C}$, a full subcategory $\mathcal{D}\subset\mathcal{C}$ is called a Serre subcategory if for every short exact sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{C}$, $Y\in\mathcal{D}$ if and only if $X,Z\in\mathcal{D}$. In particular, a Serre subcategory is closed under taking subobjects, quotients and extensions.
In a finite tensor category, every object $X$ has a socle filtration $0=X_{0}\subset X_{1}\subset...\subset X_{n}=X$ in which $X_{i+1}/X_{i}$ is the sum of simple subobjects of $X/X_{i}$. For $X\in\mathcal{C}$, we define the Loewy length to be the length of the socle filtration of $X$.
A tensor functor $F:\mathcal{C}\to\mathcal{D}$ is called surjective if every $Y\in\mathcal{D}$ is a subquotient of $F(X)$ for some $X\in\mathcal{C}$. Given a finite tensor category $\mathcal{C}$, a fiber functor is an exact faithful functor $F:\mathcal{C}\to\text{Vec}$ such that $F(\mathbbm{1})=\mathbbm{k}$ with a family of isomorphisms $\{J_{X,Y}:F(X)\otimes F(Y)\to F(X\otimes Y)|X,Y\in\mathcal{C}\}$ which satisfy the axioms of a monoidal functor.
By definition, a finite tensor category $\mathcal{D}$ has a simple extension $\mathcal{C}$ if $\mathcal{D}$ has a bimodule category which is semisimple with a unique simple object, in other words, if $\mathcal{D}$ has a fiber functor. This implies that $\mathcal{D}$ is equivalent (as a monoidal category) to a category of representations of some Hopf algebra, thus has trivial associativity constraint.
2.2. Braiding, symmetry, twists and Müger centralizer
A tensor category $\mathcal{C}$ is called braided if there exist isomorphisms $\{c_{X,Y}:X\otimes Y\to Y\otimes X\}_{X,Y\in\mathcal{C}}$ such that the following holds for objects $X,Y,Z\in\mathcal{C}$ and morphisms $\psi:X\to Z$:
(1) $c_{X\otimes Y,Z}=a_{Z,X,Y}\circ(c_{X,Z}\otimes\text{id}_{Y})\circ a^{-1}_{X,Z,Y}\circ(\text{id}_{X}\otimes c_{Y,Z})\circ a_{X,Y,Z},$
(2) $(\text{id}_{Y}\otimes\psi)\circ c_{X,Y}=c_{Z,Y}\circ(\psi\otimes\text{id}_{Y})$
for $a$ the associativity constraint of $\mathcal{C}$. For more details, see for instance [1].
For a braided finite tensor category $\mathcal{C}$, let $s$ be the squared braiding, i.e., $s_{X,Y}:=c_{Y,X}\circ c_{X,Y}$ for every $X,Y\in\mathcal{C}$. Recall that two objects $X,Y\in{\mathcal{C}}$ centralize each other if $s_{X,Y}=\text{id}_{X\otimes{Y}}$, and that the centralizer $\mathcal{D}^{\prime}$ of a full tensor subcategory $\mathcal{D}\subseteq\mathcal{C}$ is the full subcategory of $\mathcal{C}$ consisting of all objects which centralize every object of $\mathcal{D}$ (see, e.g., [11]). The notion of Müger’s centralizer enables us to distinguish between different types of braided finite tensor categories. For finite braided tensor categories $\mathcal{D}\subset\mathcal{C}$ we say that:
  • $\mathcal{D}$ is symmetric if $D\subseteq\mathcal{D}^{\prime}$.
  • $\mathcal{D}$ is Lagrangian if $\mathcal{D}=\mathcal{D}^{\prime}$.
  • $\mathcal{D}$ is central in $\mathcal{C}$ if $\mathcal{D}^{\prime}=\mathcal{C}$ (thus a central category is always symmetric).
  • $\mathcal{C}$ is non-degenerate if $\mathcal{C}^{\prime}=\text{Vec}$ (equivalently, Vec is a Lagrangian subcategory of $\mathcal{C}$).
In general, for $X,Y\in\mathcal{C}$ centralizing each other we have $s_{X,Y}=\text{id}_{X\otimes Y}\Leftrightarrow s_{Y,X}=\text{id}_{Y\otimes X}$.
By [12], we have in general for $\mathcal{D}\subset\mathcal{C}$
(3) $\text{\rm FPdim}(\mathcal{D})\text{\rm FPdim}(\mathcal{D}^{\prime})=\text{\rm FPdim}(\mathcal{C})\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$
When considering $\mathcal{D}\subset\mathcal{C}$, the centralizer of $\mathcal{D}$ in $\mathcal{D}$ is not necessarily its centralizer in $\mathcal{C}$. When needed, to avoid confusion, we will denote the centralizer of $\mathcal{D}$ in $\mathcal{D}$ by $(\mathcal{D})^{\prime}_{\mathcal{D}}$.
This also applies for three categories $\mathcal{E}\subset\mathcal{D}\subset\mathcal{C}$. Hence $(\mathcal{E})^{\prime}_{\mathcal{D}}$ might differ from $(\mathcal{E})^{\prime}_{\mathcal{C}}$ or $(\mathcal{E})^{\prime}_{\mathcal{E}}$. In all cases, the centralizer in the largest category will always be denoted without the subscript, i.e., in this case, $(\mathcal{E})^{\prime}_{\mathcal{C}}$ will simply be denoted as $\mathcal{E}^{\prime}$. Clearly, if $\mathcal{E}\subset\mathcal{D}\subset\mathcal{C}$ then $\mathcal{E}^{\prime}\subset(\mathcal{E})^{\prime}_{\mathcal{D}}\subset(\mathcal{E})^{\prime}$$_{\mathcal{E}}$.
For any $X,Y,Z\in\mathcal{C}$ we have the following equality, which is a direct calculation using the definition of braiding:
(4) $s_{X\otimes Y,Z}=a^{-1}_{X,Y,Z}\circ(\text{id}_{X}\otimes c_{Z,Y})\circ a_{X,Z,Y}\circ(s_{X,Z}\otimes\text{id}_{Y})\circ a^{-1}_{X,Z,Y}\circ(\text{id}_{X}\otimes c_{Y,Z})\circ a_{X,Y,Z}$
(5) $s_{X,Y\otimes Z}=a_{X,Y,Z}\circ(c_{Y,X}\otimes\text{id}_{Z})\circ a^{-1}_{Y,X,Z}\circ(\text{id}_{Y}\otimes s_{X,Z})\circ a_{Y,X,Z}\circ(c_{X,Y}\otimes\text{id}_{Z})\circ a^{-1}_{X,Y,Z}$
2.3. De-equivariantization
In this subsection, we briefly recall the notion of de-equivariantization and some results which we will use in the classification of braided non-semisimple near-group categories. For more details, see [1].
For a finite tensor category $\mathcal{C}$, and an algebra object $A\in\mathcal{C}$, $\textrm{Mod}_{\mathcal{C}}(A)$ is the category of right $A$-modules in $\mathcal{C}$. When $\mathcal{C}$ is braided and $A\in\mathcal{C}^{\prime}$, the category $\textrm{Mod}_{\mathcal{C}}(A)$ becomes a finite tensor category, with tensor product $\otimes_{A}$, and the free module functor $F:\mathcal{C}\to\textrm{Mod}_{\mathcal{C}}(A)$ given by $F(X)=X\otimes A$ is a surjective tensor functor.
For $X,Y\in\textrm{Mod}_{\mathcal{C}}(A)$, we denote by $\text{Hom}_{A}(X,Y)$ the vector space of $A$-module morphisms $f:X\to Y$ which are morphisms in the category $\mathcal{C}$. Notice that in general, we don’t have $\text{Hom}_{A}(X,Y)=\text{Hom}_{\mathcal{C}}(X,Y)$. For $X\in\mathcal{C}$ and $Y\in\textrm{Mod}_{\mathcal{C}}(A)$ we do have:
$\text{Hom}_{A}(X\otimes A,Y)=\text{Hom}_{\mathcal{C}}(X,\bar{Y})$
for $\bar{Y}$ the image of $Y$ under the forgetful functor $\textrm{Mod}_{\mathcal{C}}(A)\to\mathcal{C}$ (see Lemma 7.8.12 in [1]).
It is well known that for a finite group $G$, $A_{G}:=\text{Fun}(G,\mathbbm{k})$ is an algebra object in the category $\text{Rep}(G)$, see for instance [1].
Proposition 2.3.1.
Let $\mathcal{C}$ be a braided finite tensor category with braiding $c$, $G$ a finite group, and assume that $\text{Rep}(G)\subset\mathcal{C}^{\prime}$ as a symmetric subcategory, then $\textrm{Mod}_{\mathcal{C}}(A_{G})$ is a braided finite tensor category with braiding $\tilde{c}:M\otimes_{A_{G}}N\to N\otimes_{A_{G}}M$ such that the diagram
$M\otimes N$$N\otimes M$$M\otimes_{A_{G}}N$$N\otimes_{A_{G}}M$$\scriptstyle{c_{M,N}}$$\scriptstyle{\tilde{c}_{M,N}}$
commutes, and the free module functor functor $F:\mathcal{C}\to\textrm{Mod}_{\mathcal{C}}(A_{G})$ is braided.
Proof.
The proof is parallel to Proposition 8.23.8 in [1], using the inclusion as a fully faithful braided tensor functor $\text{Rep}(G)\to\mathcal{C}^{\prime}$. ∎
For a braided finite tensor category $\mathcal{C}$ for which 2.3.1 holds, we denote $\textrm{Mod}_{\mathcal{C}}(\text{Fun}(G,\mathbbm{k}))$ as $\mathcal{C}_{G}$, and call this category the $G$-de-equivariantization of $\mathcal{C}$.
Remark 2.3.2.
In the case of fusion categories, if $\mathcal{C}$ has a $G$-de-equivariantization, then we have $(\mathcal{C}_{G})^{G}\simeq\mathcal{C}$, for $\mathcal{D}^{G}$ the $G$-equivariantization of $\mathcal{D}$ by $G$ (see [1] for details on $G$-equivariantization). In finite tensor categories, this is not always true. For example, for $H$ the Sweedler’s Hopf algebra, $\text{Rep}(\mathbb{Z}_{2})$ is central in $\text{Rep}(H)$, and $\text{Rep}(H)_{\mathbb{Z}_{2}}=\text{Vec}$, so $(\text{Rep}(H)_{\mathbb{Z}_{2}})^{\mathbb{Z}_{2}}\not\simeq\text{Rep}(H)$. This happens because the fiber functor $F:\text{Rep}(H)\to\text{Vec}$ is not a braided fiber functor.
Proposition 2.3.3.
Let $\mathcal{C}$ be a braided finite tensor category with de-equivariantization $\mathcal{C}_{G}$, and $X\in\mathcal{C}$ a simple object. If $X\in\mathcal{C}$ is invertible, then $X\otimes A$ is invertible in $\mathcal{C}_{G}$.
Proof.
The free module functor is a surjective tensor functor, so by [13] it preserves Frobenius-Perron dimensions, thus $\text{\rm FPdim}_{\mathcal{C}_{G}}(X\otimes A)=1$, hence by [1], $X\otimes A$ is invertible in $\mathcal{C}_{G}$. ∎
2.4. Braided exact sequences of tensor categories
In this subsection, we recall the notion of braided exact sequences of finite tensor categories, which was introduced in [10].
A tensor functor $F:\mathcal{C}\to\mathcal{D}$ is called normal if any object $X\in\mathcal{C}$ admits a subobject $X^{\prime}\in\mathcal{C}$ such that $F(X^{\prime})$ is the largest subobject of $F(X)$ isomorphic to $(\mathbbm{1}_{\mathcal{D}})^{\oplus n}$ for $n\in\mathbb{N}$.
For $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ finite tensor categories over $\mathbbm{k}$, and tensor functors $i:\mathcal{C}_{1}\to\mathcal{C}_{2}$, $F:\mathcal{C}_{2}\to\mathcal{C}_{3}$, the sequence of tensor categories:
$\mathcal{C}_{1}\xrightarrow{i}\mathcal{C}_{2}\xrightarrow{F}\mathcal{C}_{3}$
is called an exact sequence of tensor categories if $F$ is surjective and normal, and $i$ is a full embedding such that $F\circ i(X)\in\text{Vec}$ for all $X\in\mathcal{C}_{1}$.
Remark 2.4.1.
Notice that this definition requires in particular that $\mathcal{C}_{1}$ admits a fiber functor. The notion of exact sequences was generalized to avoid this restriction by [14], but in this work, we will consider exact sequences as defined in [10].
An exact sequence of tensor categories is called braided if the categories $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ are braided and the functors $i,F$ are braided. A braided exact sequence is called modularization exact sequence if $\mathcal{C}_{3}$ is non-degenerate.
Remark 2.4.2.
If $\mathcal{E}\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}$ is a braided exact sequence, then $\mathcal{E}\subset\mathcal{C}^{\prime}$, that is, $\mathcal{E}$ lies in the Müger center of $\mathcal{C}$.
Example 2.4.3.
Assume that $\mathcal{C}$ is a braided finite tensor category, $G$ a finite group and the conditions of 2.3.1 hold, then the sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{G}$
is a braided exact sequence, for $F:\mathcal{C}\to\mathcal{C}_{G}$ the free module functor $F(X)=X\otimes A$ for $A=\text{Fun}(G,\mathbbm{k})$ the algebra object in $\text{Rep}(G)$.
Example 2.4.4.
Assume $\mathcal{C}$ is a braided pointed finite tensor category, and let $G$ be the group of invertible objects of $\mathcal{C}$. Assume $H<G$ is a maximal subgroup such that $\text{Rep}(H)\subset\mathcal{C}^{\prime}$ as a symmetric subcategory. The category $\mathcal{C}_{H}$ is a braided pointed finite tensor category which is non-degenerate (that is, $(\mathcal{C}_{H})^{\prime}=\text{Vec}$), and the group of invertible objects of $\mathcal{C}_{H}$ is $G/H$. Denoting the simple objects of $\mathcal{C}$ by $\{L_{g}\}_{g\in G}$, an invertible object $X\in\mathcal{C}_{H}$ is of the form $\oplus_{g\in B}L_{g}$ (as an object in $\mathcal{C}$) for $B$ a $H$-coset in $G$.
3. The structure of braided simple extensions
Let $\mathcal{C}$ be a braided simple extension of $\mathcal{D}$ with braiding $c$, and assume that $\mathcal{D}$ is not semisimple. Let $Q$ be the unique simple projective object in $\mathcal{C}$ which is not in $\mathcal{D}$, and let $\{L_{i}\}_{i=1}^{n}$ be the (isomorphism classes of) simple objects in $\mathcal{D}$ with projective covers $P_{i}$. The existence of the simple projective object $Q$ and the braiding on $\mathcal{C}$ enables us to understand the structure of $\mathcal{C}$ and $\mathcal{D}$.
3.1. Basic properties of simple extensions
The following lemma is true for a general simple extension $\mathcal{C}$ of $\mathcal{D}$, not necessarily braided:
Lemma 3.1.1.
Let $\mathcal{D}$ be a non semisimple finite tensor category, $\mathcal{C}$ a simple extension of $\mathcal{D}$, and let $Q$ be the simple projective object of $\mathcal{C}$. For a simple object $L\in\mathcal{D}$ with projective cover $P\in\mathcal{D}$ we have $Q\otimes Q:P]=\text{\rm FPdim}_{\mathcal{D}}(L)$.
Proof.
The category $\mathcal{D}$ has a simple extension, hence by definition has a fiber functor $\mathcal{D}\to\text{Vec}$. In particular, $\mathcal{D}$ is integral, and $\text{\rm FPdim}_{\mathcal{D}}(L)=\dim_{\mathbbm{k}}(L)$ (see [1]). $\mathcal{M}=<Q>$ is a $\mathcal{D}$-module category, and $Q$ is projective, hence we have an isomorphism $L\otimes{Q}\cong{\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}Q}$, which gives
$0\neq\oplus_{i=1}^{\dim_{\mathbbm{k}}(L)}\text{Hom}_{\mathcal{C}}(Q,Q)={\text{Hom}_{\mathcal{C}}(Q,L\otimes{Q})}=\text{Hom}_{\mathcal{C}}(Q\otimes{Q^{*}},L)=\text{Hom}_{\mathcal{C}}(Q\otimes{Q},L)$
the last equality arises from the fact that $Q$ is self-dual.
Since $L$ is simple, each such nonzero morphism is epi, so we get a morphism $P\to Q\otimes{Q}$, which implies that $P$ is a direct summand of $Q\otimes Q$, because $Q\otimes{Q}$ is projective.
By Schur’s lemma $\dim(\text{Hom}_{\mathcal{C}}(Q,Q))=1$, hence the multiplicity of $P$ in $Q\otimes{Q}$ is exactly $\dim_{\mathbbm{k}}(L)=\text{\rm FPdim}_{\mathcal{D}}(L)$. ∎
By Lemma 3.1.1, we have:
(6) $Q\otimes Q=\oplus_{i=1}^{n}\text{\rm FPdim}(L_{i})P_{i}\oplus rQ$
for some $r\in\mathbb{N}$. By definition, we have $\text{\rm FPdim}(\mathcal{C})=\text{\rm FPdim}(\mathcal{D})+\text{\rm FPdim}(Q)^{2}$, so using Eq. 6 we get that $\text{\rm FPdim}(\mathcal{C})=2\text{\rm FPdim}(\mathcal{D})+r\text{\rm FPdim}(Q)$.
3.2. Degeneracy in braided simple extensions
From here on, let $\mathcal{D}$ be a braided non-semisimple finite tensor category, $\mathcal{C}$ a braided simple extension of $\mathcal{D}$, with a unique simple projective object $Q$. Let $\mathcal{E}_{+}\subset\mathcal{D}$ be the full subcategory generated by objects $X$ which centralize $Q$.
Remark 3.2.1.
Assume we have $X,Y,Z\in\mathcal{C}$ such that $c_{X,Y}=\lambda\text{id}_{Y\otimes X}$ for some $\lambda\in\mathbbm{k}^{*}$ and $\psi:X\simeq Z$, then $c_{Z,Y}=\lambda\text{id}_{Y\otimes Z}$, by the following diagram:
$X\otimes Y$$Y\otimes X$$Z\otimes Y$$Y\otimes Z$$\scriptstyle{c_{X,Y}}$$\scriptstyle{\psi\otimes\text{id}_{Y}}$$\scriptstyle{\text{id}_{Y}\otimes\psi}$$\scriptstyle{c_{Z,Y}}$
which commutes due to the naturality of the braiding $c$. The same holds for $s_{X,Y}$ and $s_{Z,Y}$.
Proposition 3.2.2.
$\mathcal{E}_{+}$ is a symmetric tensor Serre subcategory of $\mathcal{D}$.
Proof.
For $X,Y\in\mathcal{D}$ centralizing $Q$, we have by Eq. 5 $s_{Q,X\otimes Y}=\text{id}_{Q\otimes X\otimes Y}$, so $X\otimes Y\in\mathcal{E}_{+}$. The object $Q$ is self dual, thus $\mathcal{E}_{+}$ is closed under taking duals (because $s_{Q,X}=\text{id}_{Q\otimes X}$ if and only if $s_{X,Q}=\text{id}_{X\otimes Q}$). This implies that $\mathcal{E}_{+}$ is a rigid monoidal category, and clearly it is abelian and $\mathbbm{k}$-linear as a full subcategory of $\mathcal{D}$.
Consider a short exact sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{D}$. Applying the functor $Q\otimes$- splits this sequence, because $Q$ is projective in $\mathcal{C}$ (and this is a short exact sequence in $\mathcal{C}$ as well), so we have a commutative diagram:
[…]
and the splitting of the two rows imply that $s_{Q,Y}=s_{Q,X}\oplus s_{Q,Z}$, thus $Y$ centralizes $Q$ if and only if $X$ and $Z$ do, which implies that $\mathcal{E}_{+}$ is a Serre subcategory.
Assume $X,Y$ centralize $Q$. By Eq. 4 for the triplet $(Q,X,Y)$ and $s_{Q,Y}=\text{id}_{Q\otimes Y}$ we get:
$s_{Q\otimes X,Y}=a^{-1}_{Q,X,Y}\circ(\text{id}_{Q}\otimes s_{X,Y})\circ a_{Q,X,Y},$
while on the other hand, $Q\otimes X$ is a direct sum of copies of $Q$, so by Remark 3.2.1 $s_{Q\otimes X,Y}=\text{id}_{Q\otimes X\otimes Y}$, which implies that $\text{id}_{Q}\otimes s_{X,Y}=\text{id}_{Q\otimes X\otimes Y}$, thus $s_{X,Y}=\text{id}_{X\otimes Y}$. ∎
By Corollary 8.10.8 in [1], the distinguished invertible object of $\mathcal{C}$ lies in $\mathcal{E}_{+}$.
Corollary 3.2.3.
If $Q\in\mathcal{D}^{\prime}$ then $\mathcal{D}=\mathcal{E}_{+}$ is symmetric.
Proof.
By definition, if $Q\in\mathcal{D}^{\prime}$ then $\mathcal{E}_{+}=\mathcal{D}$, and by 3.2.2, $\mathcal{D}$ is symmetric. ∎
Proposition 3.2.4.
Let $\mathcal{D}$ be a non-semisimple braided finite tensor category, $\mathcal{C}$ a braided simple extension of $\mathcal{D}$ with a simple projective object $Q$, then $s_{Q,Q}\neq\text{id}_{Q\otimes Q}$
Proof.
Assume $s_{Q,Q}=\text{id}_{Q\otimes Q}$, then by Eq. 4 for $X=Y=Z=Q$ we get $s_{Q\otimes Q,Q}=\text{id}_{Q\otimes Q\otimes Q}$. By Eq. 6, we get $s_{Q,P_{i}}=\text{id}_{Q\otimes P_{i}}$, thus $P_{i}\in\mathcal{E}_{+}$ for each $i$. By 3.2.2 $\mathcal{E}_{+}$ is a Serre subcategory, which implies that every indecomposable object $X\in\mathcal{D}$ is also in $\mathcal{E}_{+}$, thus $\mathcal{E}_{+}=\mathcal{D}$.
By 3.2.3, $\mathcal{D}$ is symmetric, and $Q$ centralizes $\mathcal{C}$ (as it centralizes $\mathcal{D}$ and $s_{Q,Q}=\text{id}_{Q\otimes Q}$), thus $\mathcal{C}$ is symmetric.
The category $\mathcal{C}$ is a symmetric finite tensor category over an algebraically closed field of characteristic 0, thus by [15], $\mathcal{C}$ is equivalent to a category of representations of some finite supergroup, but such a category has no simple projective object, unless it is semisimple itself, thus contradicting our assumption. ∎
Theorem 3.2.5.
Let $\mathcal{D}$ be a braided finite tensor category. If $\mathcal{D}$ has a braided simple extension $\mathcal{C}$, then $\mathcal{D}$ is degenerate.
Proof.
Denote $\mathcal{C}=\mathcal{D}\oplus\mathcal{M}$ and let $Q$ be the simple projective object in $\mathcal{M}$. Set $\text{\rm FPdim}(\mathcal{D})=n$ and $\text{\rm FPdim}(Q)=q$ ($q\in\mathbbm{k}^{*}$). By Eq. 6, $q^{2}=n+rq$, thus $\text{\rm FPdim}(\mathcal{C})=2n+rq$. By Eq. 3 and 3.2.2 we get in $\mathcal{C}$:
(7) $n\text{\rm FPdim}(\mathcal{D}^{\prime})=(2n+rq)\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$
which implies that $\text{\rm FPdim}(\mathcal{D}^{\prime})>\text{\rm FPdim}(\mathcal{C}^{\prime}\cap\mathcal{D})$. If $Q\notin\mathcal{D}^{\prime}$, then $\mathcal{D}^{\prime}$ (the centralizer of $\mathcal{D}$ in $\mathcal{C}$) is a full subcategory of $(\mathcal{D})^{\prime}_{\mathcal{D}}$ (the centralizer of $\mathcal{D}$ in $\mathcal{D}$), with $\text{\rm FPdim}(\mathcal{D}^{\prime})>1$, thus $\text{Vec}\subsetneq(\mathcal{D})^{\prime}_{\mathcal{D}}$, which implies that $\mathcal{D}$ is degenerate. On the other hand, if $Q\in\mathcal{D}^{\prime}$, then by 3.2.3, $\mathcal{D}$ is symmetric, hence degenerate. ∎
Proposition 3.2.6.
$\mathcal{C}^{\prime}=\mathcal{E}_{+}$. In particular, $\mathcal{E}_{+}$ is central in $\mathcal{C}$.
Proof.
We have $\mathcal{E}_{+}\subset\mathcal{D}\subset\mathcal{C}$. We will apply Eq. 3 for the pairs $(\mathcal{E}_{+},\mathcal{D})$ and $(\mathcal{E}_{+},\mathcal{C})$. We get:
(8) $\text{\rm FPdim}(\mathcal{E}_{+})\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})=\text{\rm FPdim}(\mathcal{D})\text{\rm FPdim}(\mathcal{E}_{+}\cap(\mathcal{D})_{\mathcal{D}}^{\prime}),$
(9) $\text{\rm FPdim}(\mathcal{E}_{+})\text{\rm FPdim}(\mathcal{E}_{+}^{\prime})=\text{\rm FPdim}(\mathcal{C})\text{\rm FPdim}(\mathcal{E}_{+}\cap\mathcal{C}^{\prime}).$
On one hand, we have $\mathcal{C}^{\prime}\cap\mathcal{E}_{+}\subset(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$, while on the other hand, for $X\in(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$ we have $s_{X,Y}=\text{id}_{X\otimes Y}$ for all $Y\in\mathcal{D}$ (because $X\in\mathcal{D}^{\prime}$) and for $Y=Q$ (because $X\in\mathcal{E}_{+}$), thus $X\in\mathcal{C}^{\prime}\cap\mathcal{E}_{+}$, hence $\mathcal{C}^{\prime}\cap\mathcal{E}_{+}=(\mathcal{D})_{\mathcal{D}}^{\prime}\cap\mathcal{E}_{+}$, so $\text{\rm FPdim}(\mathcal{E}_{+}\cap(\mathcal{D})_{\mathcal{D}}^{\prime})=\text{\rm FPdim}(\mathcal{E}_{+}\cap\mathcal{C}^{\prime})$. This gives the equality:
(10) $\frac{\text{\rm FPdim}(\mathcal{D})}{\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})}=\frac{\text{\rm FPdim}(\mathcal{C})}{\text{\rm FPdim}(\mathcal{E}_{+}^{\prime})}$
Let us denote $\text{\rm FPdim}(\mathcal{D})=n$, $\text{\rm FPdim}(Q)=q$, $\text{\rm FPdim}(\mathcal{E}^{\prime}_{+})=e_{c}$, $\text{\rm FPdim}((\mathcal{E}_{+})^{\prime}_{\mathcal{D}})=e_{d}$. By Eq. 6 we have $\text{\rm FPdim}(\mathcal{C})=2n+rq$. In addition, by definition $Q\in\mathcal{E}^{\prime}_{+}$, thus by calculation of Frobenius-Perron dimension we get $e_{c}=e_{d}+q^{2}=e_{d}+n+rq$. Substituting these in Eq. 10 gives:
$\frac{n}{e_{d}}=\frac{2n+rq}{e_{d}+n+rq},$
which translates to the quadratic equation:
$n^{2}+n(rq-e_{d})-e_{d}rq=0.$
This quadratic equation has two solutions: $n=e_{d}$ or n=rqn=-rq. The category $\mathcal{D}$ is integral (it has a fiber functor), and $r,q$ are non-negative [^2] Notice that $q$ is not necessarily integer, but in that case as well n=rqn=-rq is not a valid solution, because $n$ and $r$ are integers., so we must have $n=e_{d}$. This implies that:
$\text{\rm FPdim}(\mathcal{E}^{\prime}_{+})=e_{d}+n+rq=2n+rq=\text{\rm FPdim}(\mathcal{C}),$
thus $\mathcal{E}^{\prime}_{+}=\mathcal{C}$, which in turn implies that $\mathcal{E}_{+}\subseteq\mathcal{C}^{\prime}$. On the other hand, by 3.2.4, $Q\notin\mathcal{C}^{\prime}$, thus $\mathcal{C}^{\prime}\subseteq\mathcal{D}$. This implies that for any $X\in\mathcal{C}^{\prime}$, $X\in\mathcal{D}$ and $s_{Q,X}=\text{id}_{Q\otimes X}$, but this means by definition that $X\in\mathcal{E}_{+}$, thus $\mathcal{C}^{\prime}\subseteq\mathcal{E}_{+}$, completing the proof. ∎
Remark 3.2.7.
Notice that in general, we don’t have $\mathcal{E}_{+}=\mathcal{D}^{\prime}$. For instance, when $\mathcal{D}$ is pointed and $\mathcal{C}$ is non-degenerate, we will prove that $\mathcal{D}$ is slightly degenerate, thus $\mathcal{E}_{+}=\text{Vec}$ while $\mathcal{D}^{\prime}=\text{sVec}$.
4. Braided simple extensions of pointed tensor categories
In this section, we assume that $\mathcal{D}$ is a pointed braided finite tensor category, and $\mathcal{C}$ is a braided simple extension of $\mathcal{D}$. Concrete examples of such categories can be found in Appendix A in [8]. Denote the simple projective object in $\mathcal{C}$ by $Q$. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a pointed braided fusion subcategory, which we denote by $\mathcal{D}_{f}$. This category corresponds to a pair $(G,q)$ for $G$ a finite abelian group and $q$ a quadratic form on $G$. The group $G$ is simply $\text{Pic}(\mathcal{D})$ (the Picard group), and the squared braiding $s_{L_{g},L_{h}}:L_{g}\otimes L_{h}\to L_{h}\otimes L_{g}$ of two simple objects $L_{g},L_{h}\in\mathcal{D}_{f}$ is multiplication by $b(g,h)=\frac{q(gh)}{q(g)q(h)}$.
From here on, we denote the simple objects of $\mathcal{D}$ by $L_{g}$ ($g\in G$). We denote the projective cover of $L_{g}$ by $P_{g}$, and set $\bar{P}:=\oplus_{g\in\text{Pic}(\mathcal{D})}$. Clearly we have $\bar{P}\otimes L_{g}\cong\bar{P}$ for all $g\in G$.
4.1. The $Q$-braiding group of $\mathcal{C}$
Let $m_{g}$ be the order of $L_{g}$ in $\text{Pic}(\mathcal{D})$ (that is, $m_{g}$ is the minimal integer such that $L_{g}^{\otimes m_{g}}=\text{id}_{\mathcal{D}}$). By the definition of a simple extension, we have
$s_{Q,L_{g}}\in\text{Hom}_{\mathcal{C}}(Q\otimes L_{g},Q\otimes L_{g})=\text{Hom}_{\mathcal{C}}(Q,Q).$
Since $\mathbbm{k}$ is algebraically closed, we get by Schur’s lemma that $s_{Q,L_{g}}$ is simply multiplication by a scalar, which we denote by $\lambda_{g}$.
Applying Eq. 5 to the triplet $(Q,L_{g},L_{h})$, and using the fact that $s_{Q,L_{h}}$ are scalars, we get that $\lambda_{g}\lambda_{h}=\lambda_{gh}$, and $s_{Q,\mathbbm{1}}=1$. In other words, we have a group homomorphism $\psi:\text{Pic}(\mathcal{D})\to\mathbbm{k}^{*}$ given by $\psi(L_{g})=\lambda_{g}$. We define $G_{Q}:=Im(\psi)$ and call this the $Q$-braiding group of $\mathcal{D}$.
Since every $L_{g}$ has a finite order $m_{g}$, by repeatedly using Eq. 5 we see that $\lambda_{g}$ must be a root of unity, of order dividing $m_{g}$, thus $G_{Q}$ is a group of roots of unity of order dividing $|\text{Pic}(\mathcal{D})|$.
What is the kernel of the morphism $\psi$? It contains exactly all the simple objects $L_{g}\in\mathcal{D}$ such that $\lambda_{g}=1$, that is all the simple objects $L_{g}\in\mathcal{D}$ for which $s_{Q,L_{g}}=\text{id}_{Q\otimes L_{g}}$. Each such simple object lies, by definition, in $\mathcal{E}_{+}$, thus $\ker\psi=\text{Pic}(\mathcal{E}_{+})$.
In a way, the group $G_{Q}$ ”encodes” the additional braiding data of the simple extension $\mathcal{C}$ of $\mathcal{D}$.
4.2. Non-semisimple braided near-group categories
Definition 4.2.1.
Let $\mathcal{D}$ be a pointed non-semisimple finite tensor category admitting a fiber functor, and let $G=\text{Pic}(\mathcal{D})$.
  1. (1)
    A simple extension $\mathcal{C}$ of $\mathcal{D}$ is called a non-semisimple near-group category.
  2. (2)
    Denoting the unique simple projective object of $\mathcal{C}$ by $Q$, we say that $\mathcal{C}$ has a generalized fusion rule $(G,r)$ for $r$ such that Eq. 6 holds.
  3. (3)
    If $\mathcal{D}$ and $\mathcal{C}$ are braided, and the restriction of the braiding of $\mathcal{C}$ to $\mathcal{D}$ coincide with the braiding of $\mathcal{D}$, we say that $\mathcal{C}$ is a braided non-semisimple near-group category.
Proposition 4.2.2.
For a braided non-semisimple near-group category $\mathcal{C}$ with a generalized fusion rule $(G,r)$, if $r>0$ then $\mathcal{E}_{+}=\mathcal{D}$, i.e. $\mathcal{D}$ is symmetric and $\mathcal{D}^{\prime}=\mathcal{C}$.
Proof.
Consider the morphism $s_{Q\otimes Q,L_{g}}$. On one hand, by Eq. 4 for the triplet $(Q,Q,L_{g})$ we get (see also Remark 3.2.1):
$s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}=\lambda_{g}^{2}\text{id}_{\bar{P}}\oplus(\oplus_{r}\lambda_{g}^{2}\cdot\text{id}_{Q})$
We use here the isomorphism $Q\otimes L_{g}\cong Q$. On the other hand, using the naturality of the square braiding we have:
$s_{Q\otimes Q,L_{g}}=s_{\bar{P},L_{g}}\oplus(\oplus_{r}s_{Q,L_{g}})=s_{\bar{P},L_{g}}\oplus(\oplus_{r}\lambda_{g}\text{id}_{Q})$
Comparing these two expressions and using our assumption that $r>0$, we conclude that $\lambda_{g}=1$ for all $g$. This implies that $L_{g}\in\mathcal{E}_{+}$ for all $g\in G$. By 3.2.2, $\mathcal{E}_{+}$ is a Serre subcategory of $\mathcal{D}$, thus $\mathcal{E}_{+}=\mathcal{D}$, hence $\mathcal{D}$ is symmetric. By 3.2.6 we get that $\mathcal{D}^{\prime}=\mathcal{C}$. ∎
Theorem 1. Let $\mathcal{C}$ be a non-semisimple braided near-group category with generalized fusion rule $(G,r)$, then $r=0$, and consequently $\mathcal{C}$ is weakly integral.
Proof of 1.
Let $\mathcal{C}$ be a braided non-semisimple near-group category with a generalized fusion rule $(G,r)$, and assume that $r>0$. By 4.2.2, we get that $\mathcal{D}$ is a symmetric subcategory of $\mathcal{C}$ and $Q\in\mathcal{D}^{\prime}$. By [15], since $\mathcal{D}$ is not semisimple, it is braid equivalent to $\textrm{sRep}(G\ltimes W,u)$, a category of finite supermodules of the supergroup $G\ltimes W$, with $u\in G$ a central element of order $2$ acting as the parity. Since $\mathcal{D}$ is pointed, the simple objects of $\mathcal{D}$ generate a fusion subcategory, which is braid equivalent to $\text{Vec}_{G\times\mathbb{Z}/2\mathbb{Z}}$ (the central element $u$ generates $\mathbb{Z}/2\mathbb{Z}$). In particular, there exist a simple object $S\in\mathcal{D}_{f}\subset\mathcal{D}$ such that $S\otimes S=\mathbbm{1}$ and $c_{S,S}=-\tau_{S,S}$ for $\tau_{X,Y}$ the flip morphism $\tau_{X,Y}:X\otimes Y\to Y\otimes X$.
The tensor product in $\mathcal{C}$ is compatible with the left and right $\mathcal{D}$-module structure of $\mathcal{M}$, which in turn is compatible with the left and right $\mathcal{D}$-module structure of Vec, thus for $a$ the associativity constraint in $\mathcal{C}$ we have $a_{S,S,Q}=\text{id}_{S\otimes S\otimes Q}$ and $a_{Q,S,S}=\text{id}_{Q\otimes S\otimes S}$. The objects $S$ and $Q$ are simple, and $S$ is invertible, thus by abuse of notation, we can treat the morphisms $c_{S,S},c_{S,Q},c_{Q,S}$ and $a_{S,Q,S}$ as scalars.
Consider the following diagram:
$S\otimes(Q\otimes S)$$(S\otimes Q)\otimes S$$S\otimes S\otimes Q$$Q\otimes S\otimes S$$\mathbbm{1}\otimes Q$$Q\otimes\mathbbm{1}$$\scriptstyle{a^{-1}_{S,Q,S}}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}}$$\scriptstyle{\text{id}_{S}\otimes c_{S,Q}}$$\scriptstyle{c_{S\otimes S,Q}}$$\scriptstyle{c_{\mathbbm{1},Q}}$
the upper rectangle commutes by the axiom of the braiding (suppressing the associativity constraints $a_{Q,S,S}$ and $a_{S,S,Q}$ which are identities as explained above), and the lower square commutes by the naturality of the braiding. This diagram implies that $a_{S,Q,S}\cdot(c_{S,Q}^{2})=1$.
Now consider the following diagram. For ease of read, we use the notation $S^{\prime}$ to distinguish the second copy of $S$ from the first:
$S\otimes(S^{\prime}\otimes Q)$$(S\otimes S^{\prime})\otimes Q$$S\otimes(Q\otimes S^{\prime})$$(S^{\prime}\otimes S)\otimes Q$$(S\otimes Q)\otimes S^{\prime}$$S^{\prime}\otimes(S\otimes Q)$$(Q\otimes S)\otimes S^{\prime}$$(S^{\prime}\otimes Q)\otimes S$$Q\otimes(S\otimes S^{\prime})$$Q\otimes(S^{\prime}\otimes S)$$(Q\otimes S^{\prime})\otimes S$$\scriptstyle{a^{-1}_{S,S^{\prime},Q}}$$\scriptstyle{c_{S,S^{\prime}}\otimes\text{id}_{Q}}$$\scriptstyle{\text{id}_{S}\otimes c_{Q,S^{\prime}}}$$\scriptstyle{a_{S^{\prime},S,Q}}$$\scriptstyle{c_{S\otimes Q,S^{\prime}}}$$\scriptstyle{c_{S,Q}\otimes\text{id}_{S}^{\prime}}$$\scriptstyle{a_{S,Q,S^{\prime}}}$$\scriptstyle{a_{Q,S,S^{\prime}}}$$\scriptstyle{\text{id}_{S}^{\prime}\otimes c^{-1}_{Q,S}\circ a_{S^{\prime},Q,S}}$$\scriptstyle{\text{id}_{Q}\otimes c_{S,S^{\prime}}}$$\scriptstyle{a^{-1}_{Q,S^{\prime},S}}$$\scriptstyle{c_{Q,S^{\prime}}\otimes\text{id}_{S}}$
both rectangles commute by Eq. 1 for the triplet $(S,Q,S)$. From the upper rectangle, by Remark 3.2.1 for $S\otimes Q\cong Q$ and $c_{S,S^{\prime}}=-1$, we get that $a_{S,Q,S}=-1$, while the lower rectangle shows that $c_{S,Q}=c_{Q,S}$.
$a_{S,Q,S}=-1$ implies that $(c_{S,Q})^{2}=-1$, that is $c_{Q,S}^{2}=-1$, but $c_{Q,S}\cdot{c_{S,Q}}=1$ (because $S\in\mathcal{E}_{+}$), thus $c_{Q,S}=-c_{S,Q}$, which yields a contradiction.
By Eq. 6, if $r=0$ then $Q\otimes Q=\oplus_{g\in G}P_{g}$ for $P_{g}$ the projective cover of the simple object $L_{g}$, and all the projective covers of simple objects have the same Frobenius-Perron dimension, thus $\text{\rm FPdim}(\mathcal{C})=2|G|\text{\rm FPdim}(P_{e})$ for $P_{e}$ the projective cover of $\mathbbm{1}$, hence $\mathcal{C}$ is weakly integral. ∎
As a result of 1, when we consider non-semisimple braided near-group categories $\mathcal{C}$, the generalized fusion rule is always $(\text{Pic}(\mathcal{D}),0)$, so as an abbreviation, we will simply say that $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category.
4.3. De-equivariantization of near-group categories
In this subsection, we prove that every $\text{Pic}(\mathcal{D})$-braided near-group category is ”an extension” of a non-degenerate braided non-semisimple near-group category by some symmetric fusion category, that is, there is a braided exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{nd}$
for $\mathcal{C}_{nd}$ a non-degenerate braided non-semisimple near-group category. This implies that the classification of braided non-semisimple near-group categories restricts to the classification of non-degenerate braided non-semisimple near-group categories, in a way.
Proposition 4.3.1.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category for some $\mathcal{D}$. Assume that for $H<\text{Pic}(\mathcal{D})$ we have $\text{Rep}(H)\subset\mathcal{C}^{\prime}$, then $\mathcal{C}_{H}$ is a braided non-semisimple near-group category.
Proof.
By 2.3.1 $\mathcal{C}_{H}$ is a well-defined braided finite tensor category, and the free module functor $F:\mathcal{C}\to\mathcal{C}_{H}$ is a braided tensor functor. This functor is surjective and normal by [10].
Let $Y\in\mathcal{C}_{H}$ be a simple object. By Lemma 2.1 in [9], $Y$ is a quotient of $F(X)$ for some $X\in\mathcal{C}$, that is we have a surjective morphism $\psi:F(X)\to Y$. Since $Y$ is simple and $F$ is a tensor functor (hence additive), we may assume that $X\in\mathcal{C}$ is indecomposable, and we can choose $X$ such that $\psi:F(X)\to Y$ is non-zero and $X$ has minimal Loewy length. Let $L$ be the top of $X$, then $L=Q$ or $L\in\mathcal{D}$, and we have a surjective morphism $\varphi:X\to L$.
Assume $L\in\mathcal{D}$. We have an exact sequence in $\mathcal{D}$:
$0\to Z\to X\xrightarrow{\varphi}L\to 0$
and $\text{\rm FPdim}(L)=1$, thus by 2.3.3, $F(L)$ is also invertible, and $F(\varphi):F(X)\to F(L)$ is a surjective morphism (because $F$ is exact). If $F(L)\simeq Y$ we are done, otherwise $\varphi|_{F(Z)}:F(Z)\to Y$ must be non-zero, but the Loewy length of $Z$ is smaller than that of $X$, which contradicts the choice of $X$.
If $L=Q$, by the assumption that $X$ is indecomposable, we must have $X=Q$. By [13], surjective tensor functors map projective objects to projective objects, thus $F(X)=Q\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$ for $A:=\text{Fun}(H,\mathbbm{k})$ the algebra object in $\text{Rep}(H)$. By Lemma 7.8.12 in [1], $\text{Hom}_{A}(Q\otimes A,Y)=\text{Hom}_{\mathcal{C}}(Q,\bar{Y})$ for $\bar{Y}$ the $A$-module $Y$ considered as an object in $\mathcal{C}$, thus as an object in $\mathcal{C}$, we have $Y=Z\oplus Q^{\oplus m}$ for some $m\in\mathbb{N}$, $Z\in\mathcal{D}$.
The category $\textrm{Mod}_{\mathcal{C}}(A)$ is additive, and as an object in $\mathcal{C}$, $A$ is a direct sum of simple objects in $\mathcal{D}$, thus $Q\otimes A$ is a direct sum of copies of $Q$, while $Z\otimes A\in\mathcal{D}$. Since we assume that $Y$ is simple as an $A$-module, this implies that $Z=0$.
The algebra $A$ is separable (as an algebra object in $\text{Rep}(H)$), so $Y$ is a direct summand of $Y\otimes A$ in $\textrm{Mod}_{\mathcal{C}}(A)$, so to prove that $Y$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, it is enough to prove that $Y\otimes A$ is projective in $\textrm{Mod}_{\mathcal{C}}(A)$, but this is clear because of Lemma 7.8.12 in [1] and the fact that $\bar{Y}$ is a direct sum of copies of $Q$. In addition, it is clear from the proof that such $Y$ is uniquely determined by the fact that $L=Q$.
We conclude that $\textrm{Mod}_{\mathcal{C}}(A)$ is a braided, with a unique simple projective object, and all the other objects are invertible, hence it is a braided non-semisimple near-group category. ∎
Corollary 4.3.2.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group, $H<\text{Pic}(\mathcal{D})$ such that $\text{Rep}(H)\subset\mathcal{C}^{\prime}$, and for $X\in\mathcal{C}_{H}$ denote by $\bar{X}$ the underlying object in $\mathcal{C}$.
  1. (1)
    If $Y\in\mathcal{C}_{H}$ is the unique simple projective object, then $\bar{Y}=Q^{\oplus m}$ for $m\in\mathbb{N}$.
  2. (2)
    If $Y\in\mathcal{C}_{H}$ is invertible, then $\bar{Y}$ is a direct sum of invertible objects in $\mathcal{D}$.
Proof.
The first claim follows directly from the proof of 4.3.1. The second claim follows from Example 2.4.4. ∎
Proposition 4.3.3.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category, then there exist $H<\text{Pic}(\mathcal{D})$ such that $\mathcal{C}_{H}$ is a non-degenerate braided near-group category.
Proof.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, $\mathcal{E}_{+}$ is a central symmetric subcategory of $\mathcal{C}$ and $\mathcal{D}$, thus in particular it is a symmetric pointed tensor category. Set $H:=\text{Pic}(\mathcal{E}_{+})$, then $\text{Rep}(H)$ is also symmetric and central in $\mathcal{D}$ and $\mathcal{C}$. By 4.3.1, $\mathcal{C}_{H}$ is a braided non-semisimple near-group category, so we need only to prove that $\mathcal{C}_{H}$ is non-degenerate. Set $A:=\text{Fun}(H,\mathbbm{k})$.
Let $Y\in\mathcal{C}_{H}$ be the unique simple projective object, and let $L\in\mathcal{C}_{H}$ be a simple invertible object. By 4.3.2, we have $\bar{Y}=Q^{\oplus m}$ for some $m\in\mathbb{N}$ and $\bar{L}=\oplus_{f\in B}L_{f}$ for $B$ a coset of $H$ in $G$. Assume that $\tilde{s}_{Y,L}=\text{id}_{Y\otimes_{A}L}$ for $\tilde{s}$ the squared braiding in $\mathcal{C}_{H}$. By 2.3.1 we have a commutative diagram:
$\bar{Y}\otimes\bar{L}$$\bar{L}\otimes\bar{Y}$$Y\otimes_{A}L$$L\otimes_{A}Y$$\scriptstyle{s_{\bar{Y},\bar{L}}}$$\scriptstyle{\tilde{s}_{Y,L}}$
thus $s_{\bar{Y},\bar{L}}=\text{id}_{\bar{Y}\otimes\bar{L}}$. By the definition of $\mathcal{E}_{+}$, since $\bar{Y}=Q^{\oplus m}$, we have $L_{f}\in\mathcal{E}_{+}$ for all $f\in B$. This implies (by Example 2.4.4) that $B$ is the trivial coset, that is $\bar{L}=\oplus_{h\in H}L_{h}$, hence $L=\mathbbm{1}_{\mathcal{C}_{H}}$, so in $\mathcal{C}_{H}$ the object $Y$ is centralized only by $\mathbbm{1}_{\mathcal{C}_{H}}$, thus $\mathcal{C}_{H}$ is non-degenerate as required. ∎
Theorem 2. Let $\mathcal{C}$ be a braided non-semisimple near-group category, then there exist a non-degenerate braided non-semisimple near-group category $\mathcal{D}$ and a finite group $G$ such that $\mathcal{C}$ fits in a modularization exact sequence:
$\text{Rep}(G)\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{D}.$
Proof of 2.
Let $\mathcal{C}$ be a braided non-semisimple near-group category, i.e. a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$. By 1, $\mathcal{C}$ is a $\text{Pic}(\mathcal{D})$-braided near-group category. By 3.2.2 and 3.2.6, the category $\mathcal{E}_{+}$ generated by objects in $\mathcal{D}$ that centralize the unique simple projective object $Q\in\mathcal{C}$ is a symmetric subcategory of $\mathcal{D}$ which is central in $\mathcal{C}$. Hence, by 4.3.3, taking $H:=\text{Pic}(\mathcal{E}_{+})$ we get that $\mathcal{C}_{H}$ is a non-degenerate $G_{Q}$-braided near-group category and we get a modularization exact sequence:
$\text{Rep}(\text{Pic}(\mathcal{E}_{+}))\xrightarrow{i}\mathcal{C}\xrightarrow{F}\mathcal{C}_{H}.$
4.4. Non-degenerate near-group categories
As a result of 2, in order to fully classify braided non-semisimple near-group categories, we need to classify non-degenerate braided non-semisimple near-group categories. This section is devoted to this classification.
Proposition 4.4.1.
If $\mathcal{C}$ is a non-degenerate braided non-semisimple near-group category with a generalized fusion rule $(G,0)$, then $G=\mathbb{Z}_{2}$.
Proof.
By 3.2.2, since $\mathcal{C}$ is non-degenerate we have $\text{\rm FPdim}(\mathcal{E}_{+})=1$, thus by Eq. 7 (together with $r=0$) we get that $\text{\rm FPdim}(\mathcal{D}^{\prime})=2$, that is $\mathcal{D}$ is slightly degenerate.
Let $L_{g}\in\mathcal{D}$ be a non-trivial simple object, and denote $s_{Q,L_{g}}=\lambda_{g}$. By Eq. 4 for the triplet $(Q,Q,L_{g})$ we get that $s_{Q\otimes Q,L_{g}}=\lambda_{g}^{2}\text{id}_{Q\otimes Q\otimes L_{g}}$. On the other hand, by Eq. 6, the naturality of the braiding, and the fact that $s_{Q\otimes Q,L_{g}}$ is a scalar map, we get that $s_{\bar{P},L_{g}}=\lambda_{g}^{2}\bar{P}\otimes L_{g}$. In particular, for $P_{e}$ the projective cover of $\mathbbm{1}$ we have $s_{P_{e},L_{g}}=\lambda_{g}^{2}\text{id}_{P_{e}\otimes L_{g}}$. $\mathcal{C}$ is non-degenerate, thus unimodular (by [1]), so $\mathbbm{1}$ is the socle of $P_{e}$, thus $\lambda^{2}_{g}\text{id}_{L_{g}}=s_{\mathbbm{1},L_{g}}=\text{id}_{L_{g}}$, hence $\lambda_{g}^{2}=1$.
We conclude that for each projective cover $P_{h}$ of $L_{h}\in\text{Pic}(\mathcal{D})$ we have $s_{P_{h},L_{g}}=\text{id}_{P_{h}\otimes L_{g}}$, thus for each projective object $P\in\mathcal{D}$ we must have $s_{P,L_{g}}=\text{id}_{P\otimes L_{g}}$. Let $X\in\mathcal{D}$ be indecomposable, then $X$ has a projective cover $P^{\prime}$, a surjective map $\pi:P^{\prime}\to X$, and we get a commutative diagram:
$P^{\prime}\otimes L_{g}$$P^{\prime}\otimes L_{g}$$X\otimes L_{g}$$X\otimes L_{g}$$\scriptstyle{s_{P^{\prime}\otimes L_{g}}}$$\scriptstyle{\pi\otimes\text{id}_{L_{g}}}$$\scriptstyle{\pi\otimes\text{id}_{L_{g}}}$$\scriptstyle{s_{X,L_{g}}}$
which implies that $s_{X,L_{g}}=\text{id}_{X\otimes L_{g}}$ as well. Overall, we get that for $L_{g}$ a non-trivial simple object in $\mathcal{D}$, $L_{g}\in\mathcal{D}^{\prime}$, but $\mathcal{D}^{\prime}$ is slightly degenerate, thus $L_{g}$ must be the unique non-trivial invertible element in $\mathcal{D}^{\prime}\simeq\text{sVec}$, hence $G=\mathbb{Z}_{2}$ as required.
Theorem 3. Every non-degenerate non-semisimple braided near-group category $C$ is a braided simple extension of a category $\mathcal{D}$ braid equivalent to $\textrm{sRep}(W\oplus W^{*})$, with $\textrm{sRep}(W)$ a Lagrangian subcategory, for some purely odd supervector space $W$.
Proof of 3.
Let $\mathcal{C}$ be a non-degenerate, braided non-semisimple near-group category with fusion rule $(G,0)$. By 4.4.1, we have $G=\mathbb{Z}_{2}$. This implies that $\mathcal{D}$ is a braided finite tensor category with only two simple objects, both invertible. Since $\mathcal{D}_{f}=\text{sVec}$ is a symmetric subcategory of $\mathcal{D}$, there is some Lagrangian subcategory $\mathcal{D}_{0}$ in $\mathcal{D}$, and we must have $\text{sVec}\subset\mathcal{D}_{0}$, so by [15], $\mathcal{D}_{0}$ is equivalent to a category of representations of some supergroup, and since $|G|=2$ this supergroup must be $\wedge W$ for some purely odd supervector space $W$, that is $\mathcal{D}_{0}\cong\textrm{sRep}(W)$. Now using [8] we get the required result. ∎
In 1 we proved that every braided non-semisimple near-group category is weakly integral. Using 2 and 3 we get a stronger result:
Corollary 4.4.2.
A braided non-semisimple near-group category $\mathcal{C}$ is non-integral weakly integral.
Proof.
by 1 we know that $\mathcal{C}$ is weakly integral, so we need to prove it is not integral. By 2, we have an exact sequence:
$0\to\text{Rep}(G)\to\mathcal{C}\to\bar{\mathcal{C}}\to 0$
for $\bar{\mathcal{C}}$ a non-degenerate braided non-semisimple near-group category. By 3, $\bar{\mathcal{C}}$ is a simple extension of $\bar{\mathcal{D}}$ which is braid equivalent to $\textrm{sRep}(W\oplus W^{*})$. A direct calculation shows that $\text{\rm FPdim}(\bar{\mathcal{D}})=2^{2\dim(W)+1}$, so denoting by $Q_{0}$ the unique simple projective object of $\bar{\mathcal{C}}$, we have $\text{\rm FPdim}(Q_{0})=\sqrt{2}\cdot 2^{\dim(W)}$. Denoting by $Q$ the unique simple object in $\mathcal{C}$, we know that the functor $\mathcal{C}\to\bar{\mathcal{C}}$ is the free module functor and it maps $Q$ to $Q_{0}$. Since this is a surjective functor, by [13] it preserve Frobenius-Perron dimensions, thus $\text{\rm FPdim}(Q)=\sqrt{2}\cdot^{\dim(W)}\notin\mathbb{N}$. ∎
Using 2 and 3 we get the following result:
Proposition 4.4.3.
Let $\mathcal{C}$ be a $\text{Pic}(\mathcal{D})$-braided non-semisimple near-group category, then $\mathcal{D}$ has a simple object $S$ such that $S\otimes S=\mathbbm{1}$ and $S\in\mathcal{D}^{\prime}$.
Proof.
By 2 and 3, for $\mathcal{E}_{+}=\mathcal{C}^{\prime}$ we have a modularization exact sequence
$\text{Rep}(\text{Pic}(\mathcal{E}_{+}))\to\mathcal{C}\xrightarrow{F}\mathcal{C}_{\text{Pic}(\mathcal{E}_{+})}$
and $\mathcal{C}_{\text{Pic}(\mathcal{E}_{+})}$ is a non-degenerate braided non-semisimple near-group category. Denote $G:=\text{Pic}(\mathcal{C})$, $H:=\text{Pic}(\mathcal{E}_{+})$ (so $H<G)$, and denote the simple invertible objects of $\mathcal{C}$ by $L_{g}$ for $g\in G$.
$\mathcal{C}_{H}$ has a unique non-trivial invertible object $\tilde{S}$. The functor $F$ is dominant, so we have $X\in\mathcal{C}$ such that $\tilde{S}$ is a subobject of $F(X)$. Let $L_{g}$ be the socle of $F(X)$, then $F(L_{g})\simeq\tilde{S}$, hence $g\notin H$.
The functor $F$ is a tensor functor, so $F(L_{g}\otimes L_{g})\simeq F(L_{g})\otimes F(L_{g})\simeq\tilde{S}\otimes\tilde{S}=\mathbbm{1}$, thus $g^{2}\in H$.
$G$ is an abelian group, so $(gH)^{2}=g^{2}H=H$, that is, $H$ is a subgroup of index $2$, thus there is $g_{0}\in G\setminus H$ such that $g_{0}^{2}=1_{G}$. Set $S:=L_{g_{0}}$, then $S\otimes S=\mathbbm{1}_{\mathcal{C}}$.
By 3, the category $\mathcal{C}_{H}$ has a braided subcategory $\tilde{\mathcal{D}}$ which is tensor equivalent to $\textrm{sRep}(W\oplus W^{*})$ for some purely odd supervector space $W$, and by [8], $\tilde{S}\in\tilde{\mathcal{D}}^{\prime}$.
By the proof of 4.3.1, the functor $F:\mathcal{C}\to\mathcal{C}_{H}$ restricts (as a braided tensor functor) to $F_{\mathcal{D}}:\mathcal{D}\to\tilde{\mathcal{D}}$, thus by 2.3.1, we have $s_{X,S}=\text{id}_{X\otimes S}$ for all $X\in\text{Pic}(\mathcal{D})$. ∎
This proposition, along with the previous results, gives the following:
Theorem 5. Every braided non-semisimple near-group category $\mathcal{C}$ corresponds to a pair $(G,n_{p})$ such that:
  1. (1)
    $G$ is a group of order $2^{n_{g}}$ with a central element of order $2$.
  2. (2)
    $n_{p}\in\mathbb{N}$ and for $P_{\mathbbm{1}}$ the projective cover of $\mathbbm{1}_{\mathcal{C}}$, $\dim(P_{e})=2^{n_{p}}$ as a vector space.
  3. (3)
    $n_{p}+n_{g}$ is an odd number.
Proof.
By definition, $\mathcal{C}$ is a braided simple extension of a braided pointed finite tensor category $\mathcal{D}$, and denote $G:=\text{Pic}(\mathcal{D})$. By 4.4.3, $G$ has a central element of order $2$. By the proof of 4.4.2, there exist $l\in\mathbb{N}$ such that $\text{\rm FPdim}(Q)=\sqrt{2}\cdot 2^{l}$. By 1 and Eq. 6 we get that $\text{\rm FPdim}(\mathcal{D})=2^{2l+1}$. On the other hand, $\mathcal{D}$ is pointed, so denoting its simple objects by $L_{g}$ and their projective covers by $P_{g}$, for $g\in G$, we know that $\text{\rm FPdim}(P_{g})$ is invariant to $g$, and $\text{\rm FPdim}(\mathcal{D})=|G|\cdot\text{\rm FPdim}(P_{e})$ (for $e\in G$ the unit). This imply that there exist $n_{g}\in\mathbb{N}$ such that $|G|=2^{n_{g}}$ and $\text{\rm FPdim}(P_{e})=2^{2l+1-n_{g}}$, that is $n_{p}=2l+1-n_{g}$. Since $\mathcal{D}$ is integral, $\dim(P_{e})=\text{\rm FPdim}(P_{e})=2^{n_{p}}$, and $n_{p}+n_{g}=2l+1$ is an odd number. ∎